### v2

# Crepant Transformation correspondence For Toric Stack Bundles

Abstract.
We prove a crepant transformation correspondence in genus zero Gromov-Witten theory for toric stack bundles related by crepant wall-crossings of the toric fibers. Specifically, we construct a symplectic transformation that identifies $I$-functions of toric stack bundles suitably analytically continued using Mellin-Barnes type integrals. We compare our symplectic transformation with a Fourier-Mukai isomorphism between the $K$-groups.
0. Introduction
0.1. Result
In toric geometry, there is a natural way to construct birational maps between toric stacks by presenting toric stacks as GIT quotients (see Section 2.2 for a review) and consider variations of GIT quotients. Certain such VGIT yield crepant birational maps (also known as K-equivalence).
More precisely, let $X_{\pm}$ be toric Deligne-Mumford stacks with $\varphi:X_{+}\dashrightarrow X_{-}$ the birational map arising as the stability parameter crosses a wall. Under certain assumption (see Section 2.3), there is a common toric blow-up $\tilde{X}$ with projective birational morphisms $f_{\pm}:\tilde{X}\rightarrow X_{\pm}$ such that $f_{-}=\varphi\circ f_{+}$ and $f^{\star}_{+}K_{X_{+}}=f^{\star}_{-}K_{X_{-}}$, i.e. $X_{+}$ and $X_{-}$ are $K$-equivalent. In [8], Coates, Iritani and Jiang considered this situation and proved a crepant transformation correspondence for genus 0 descendant Gromov-Witten theory of $X_{\pm}$. Moreover, they relate this correspondence in Gromov-Witten theory with Fourier-Mukai transform between equivariant derived categories of $X_{\pm}$ studied in [9].
This paper establishes a correspondence in genus 0 Gromov-Witten theory for crepant birational maps of toric stack bundles arising from toric wall-crossings. As described in Section 2.4.1 below, we consider toric stack bundles $\mathcal{P}_{\pm}$ with toric fiber $X_{\pm}$ related by a crepant toric wall-crossing. We have toric stack bundles $\tilde{\mathcal{P}}$ with fiber $\tilde{X}$ and $\mathcal{P}_{0}$ with fiber $X_{0}$, where $\tilde{X}$ is the common toric blow-up of $X_{\pm}$ and $X_{0}$ is the common blow-down. We have the following diagram with map $\varphi$ induced from the toric wall-crossing on the fiber, $f_{\pm}$ and $g_{\pm}$ induced from the toric blow-ups and blow-downs on the fiber:
$\tilde{\mathcal{P}}$$\mathcal{P}_{+}$$\mathcal{P}_{-}$$\mathcal{P}_{0}.$$\scriptstyle{f_{+}}$$\scriptstyle{f_{-}}$$\scriptstyle{\varphi}$$\scriptstyle{g_{+}}$$\scriptstyle{g_{-}}$
The main result of this paper is the following:
Theorem 0.1.
Let $\tilde{\mathcal{H}}(\mathcal{P}_{\pm})$ be multi-valued version of Givental’s symplectic vector space for $\mathcal{P}_{\pm}$. There exist a degree-preserving $R_{T}(\!(z^{-1})\!)$-linear symplectic transformation $\mathbb{U}:\tilde{\mathcal{H}}(\mathcal{P}_{-})\rightarrow\tilde{\mathcal{H}}(\mathcal{P}_{+})$ such that:
  1. (1)
    $\mathbb{U}$ identifies the $I$-functions of $\mathcal{P}_{\pm}$: $I_{+}(y,z)=\mathbb{U}I_{-}(y,z)$ after analytic continuation;
  2. (2)
    $\mathbb{U}\circ(g^{\star}_{-}v\cup)=(g^{\star}_{+}v\cup)\circ\mathbb{U}$ for all $v\in H^{2}_{T}(\mathcal{P}_{0})$;
  3. (3)
    there exists a Fourier-Mukai transformation $\mathbb{FM}:K^{0}_{T}(\mathcal{P}_{-})\rightarrow K^{0}_{T}(\mathcal{P}_{+})$ such that the following diagram commutes:
    (0.1) $K^{0}_{T}(\mathcal{P}_{-})$$K^{0}_{T}(\mathcal{P}_{+})$$\tilde{\mathcal{H}}(\mathcal{P}_{-})$$\tilde{\mathcal{H}}(\mathcal{P}_{+}).$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{{\Psi}_{-}}$$\scriptstyle{{\Psi}_{+}}$$\scriptstyle{\mathbb{U}}$
Here $\Psi_{\pm}=\tilde{\Psi}_{\pm}\circ\tilde{\text{ch}}$} and $\tilde{\Psi}_{\pm}$ are defined in (3.8) below.
By [14], the $I$-functions $I_{\pm}$ of $\mathcal{P}_{\pm}$, whose definition is recalled in (3.1), lie on the Lagrangian cones of genus 0 Gromov-Witten theory of $\mathcal{P}_{\pm}$. Therefore, Theorem 0.1 can be viewed as a crepant transformation correspondence for genus 0 Gromov-Witten theory of $\mathcal{P}_{\pm}$.
Our approach to Theorem 0.1 is an adaptation of the method used in [8] for the toric case, replacing the toric mirror theorem of [7] by its toric stack bundles counterpart in [14]. In fact, [8] is our main motivation for considering Theorem 0.1 and has been the main source of guidance in our study.
The rest of this paper is organized as follows. In Section 1 we briefly introduce Gromov-Witten theory for Deligne-Mumdford stacks and Givental’s symplectic formalism. In Section 2 we recall the definitions and basic results for toric stacks and toric stack bundles, and discuss their toric wall-crossings. In Section 3 we prove Theorem 0.1.
0.2. Acknowledgment
We thank the referee for helpful comments and suggestions. H.-H. T. is supported in part by Simons foundation collaboration grant.
1. Preliminaries on Gromov-Witten Invariants
This Section is a collection of basic facts about Gromov-Witten theory. The purpose of this Section is mainly to set up notations we need. Our presentation closely follows those in related papers including [8].
1.1. Chen-Ruan Orbifold Cohomology
Let $X$ be a smooth proper Deligne-Mumford stack of finite type over $\mathbb{C}$. Let $\lvert X\rvert$ be its coarse moduli space and $IX=X\times_{\lvert X\rvert}X$ the inertia stack of $X$. The stack $IX$ admits a decomposition into connected components
$IX=\bigsqcup_{v\in\mathsf{B}}X_{v},$
with a “main component” $X_{0}\simeq X$. There is a involution map $\text{inv}:IX\rightarrow IX$ which restricts to an isomorphism between corresponding components $\text{inv}_{i}:X_{i}\rightarrow X_{\text{inv}(i)}$, and becomes identity on $X_{0}$.
The Chen-Ruan orbifold cohomology $H^{\bullet}_{\text{CR}}(X)$ of $X$ is the cohomology of the inertia stack $IX$ with grading shifted by a rational number $\iota_{v}$ called age:
$H^{k}_{\text{CR}}(X)=\bigoplus_{v\in\mathsf{B}:k-2\iota_{v}\in 2\mathbb{Z}}H^{k-2\iota_{v}}(X_{v};\mathbb{C})$
$H^{\bullet}_{\text{CR}}(X)$ carries the orbifold Poincaré pairing: for $\alpha,\beta\in H^{\bullet}_{\text{CR}}(X)$, define
$(\alpha,\beta)_{\text{CR}}=\int_{IX}\alpha\cup\text{inv}^{*}\beta.$
Assume $\alpha\in H^{\bullet}(X_{i})$ and $\alpha\in H^{\bullet}(X_{j})$, then $(\alpha,\beta)_{\text{CR}}$ is only non-zero when $j=\text{inv}(i)$, in which case
$(\alpha,\beta)_{\text{CR}}=\int_{X_{i}}\alpha\cup\text{inv}^{*}\beta$
If $X$ admits [^1] When $X$ admits a $T$-action, we can relax the properness requirement: we only assume that the $T$-fixed locus $X^{T}\subset X$ is proper. an action of an algebraic torus $T\cong(\mathbb{C}^{\times})^{m}$, we can define the $T$-equivariant Chen-Ruan orbifold cohomology of $X$, $H^{\bullet}_{\text{CR},T}(X)$, to be the $T$-equivariant cohomology of the inertia stack with age shifting:
$H^{k}_{\text{CR},T}(X)=\bigoplus_{v\in\mathsf{B}:k-2\iota_{v}\in 2\mathbb{Z}}H^{k-2\iota_{v}}_{T}(X_{v};\mathbb{C}).$
$H^{\bullet}_{\text{CR},T}(X)$ carries the $T$-equivariant orbifold Poincaré pairing: for $\alpha,\beta\in H^{\bullet}_{\text{CR},T}(X)$, define
$(\alpha,\beta)_{\text{CR}}=\int_{IX}\alpha\cup\text{inv}^{*}\beta\in S_{T}$
using the Atiyah-Bott localization formula. Here $S_{T}$ is the localization of $R_{T}=H^{\bullet}_{T}(\text{pt};\mathbb{C})$ with respect to the set of non-zero homogeneous elements.
1.2. Gromov-Witten Invariant and Quantum Cohomology
[…]
Given $\alpha_{1},\dots,\alpha_{n}\in H^{\bullet}_{\text{CR},T}(X)$ and non-negative integers $k_{1},\dots,k_{n}$, we have the associated $T$-equivariant Gromov-Witten invariant, defined [^2] If $X$ is not proper but $X^{T}$ is proper, this is defined using localization formula for virtual fundamental classes. to be
$\langle\alpha_{1}\psi^{k_{1}},\dots,\alpha_{n}\psi^{k_{n}}\rangle^{X}_{g,n,d}=\int_{[X_{g,n,d}]^{\text{vir}}}\prod_{i=1}^{n}(\text{ev}^{*}_{i}\alpha_{i})\psi_{i}^{k_{i}}\in S_{T}.$
[…]
$\text{NE}(X)_{\mathbb{Z}}=\{d\in H_{2}(\lvert X\rvert;\mathbb{Z}):d\in\text{NE}(X)\}.$
Denote by $Q$ the Novikov variable, so for a ring $R$, we have
$R\llbracket Q\rrbracket=\left\{\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}a_{d}Q^{d}:a_{d}\in R\right\}.$
For $\mathbf{t}=\mathbf{t}(z)=t_{0}+t_{1}z+t_{2}z^{2}+\cdots\in H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}S_{T}[z]\llbracket Q\rrbracket$, define
$\langle\mathbf{t},\dots,\mathbf{t}\rangle^{X}_{g,n,d}=\langle\mathbf{t}(\psi),\dots,\mathbf{t}(\psi)\rangle^{X}_{g,n,d}=\sum_{k_{1},\dots,k_{n}\geq 0}\langle t_{k_{1}}\psi^{k_{1}},\dots,t_{k_{n}}\psi^{k_{n}}\rangle^{X}_{g,n,d}.$
The genus-$g$ descendant potential of $X$ is defined to be
$\mathcal{F}^{g}_{X}(\mathbf{t})=\sum_{n\geq 0}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\frac{Q^{d}}{n!}\langle\mathbf{t},\dots,\mathbf{t}\rangle^{X}_{g,n,d}.$
1.3. Equivariant Quantum Cohomology
Let $\{\phi_{i}\}$ be a basis of $H_{\text{CR}}^{\bullet}(X)$. Let $\eta_{ij}=(\phi_{i},\phi_{j})_{\text{CR}}$ and $\eta^{ij}$ be the entries of inverse matrix of $(\eta_{ij})$. For $\tau\in H_{\text{CR}}^{\bullet}(X)$, we define the quantum product at $\tau$ by
$\phi_{i}\bullet_{\tau}\phi_{j}$ $=\sum_{k,l}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{n=0}^{\infty}\frac{Q^{d}}{n!}\langle\phi_{i},\phi_{j},\phi_{k},\tau,\dots,\tau\rangle_{0,n+3,d}^{X}\eta^{kl}\phi_{l}.$
Equivalently,
$(\phi_{i}\bullet_{\tau}\phi_{j},\phi_{k})=\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{n=0}^{\infty}\frac{Q^{d}}{n!}\langle\phi_{i},\phi_{j},\phi_{k},\tau,\dots,\tau\rangle_{0,n+3,d}^{X}.$
For equivariant orbifold cohomology $H^{\bullet}_{\text{CR},T}(X)$ and $\tau\in H^{\bullet}_{\text{CR},T}(X)$, we define the quantum product analogously, so
$(\phi_{i}\bullet_{\tau}\phi_{j},\phi_{k})_{\text{CR}}=\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{n=0}^{\infty}\frac{Q^{d}}{n!}\langle\phi_{i},\phi_{j},\phi_{k},\tau,\dots,\tau\rangle_{0,n+3,d}^{X},$
or
$\phi_{i}\bullet_{\tau}\phi_{j}$ $=\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{n=0}^{\infty}\frac{Q^{d}}{n!}\text{inv}^{*}\text{ev}_{3,*}\left(\text{ev}^{*}_{1}(\phi_{i})\text{ev}^{*}_{2}(\phi_{j})\prod_{l=4}^{n+3}\text{ev}^{*}_{l}(\tau)\cap[X_{0,n+3,d}]^{\text{vir}}\right).$
The quantum product $\bullet_{\tau}$ defines an associative and commutative ring structure on $H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}R_{T}[z]\llbracket\tau,Q\rrbracket$, which is called the equivariant quantum cohomology of $X$.
1.4. Givental’s Symplectic Formalism
The original reference for Givental’s formalism is [10]. See [6] for a presentation of its orbifold version.
Givental’s symplectic space is defined to be
$\mathcal{H}=H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}S_{T}(\!(z^{-1})\!)\llbracket Q\rrbracket,$
where $S_{T}(\!(z^{-1})\!)$ is the ring of formal Laurent series in $z^{-1}$ with coefficients in $S_{T}$. The symplectic form on the space is given by
$\Omega(f,g)=-\text{Res}_{z=\infty}(f(-z),g(z))_{\text{CR}}dz,$
for any $f,g\in\mathcal{H}$, i.e. the coefficient of $z^{-1}$ of the orbifold Poincaré pairing of f(z)f(-z) and $g(z)$. This space has a standard polarization
$\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-},$
where
$\mathcal{H}_{+}=H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}S_{T}[z]\llbracket Q\rrbracket,$
$\mathcal{H}_{-}=z^{-1}H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}S_{T}\llbracket z^{-1}\rrbracket\llbracket Q\rrbracket.$
The subspaces $\mathcal{H}_{\pm}\subset\mathcal{H}$ are the maximal isotropic subspaces with respect to $\Omega$, and $\Omega$ induces a non-degenerate pairing between $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$. This identifies $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ with the total space of the cotangent bundle $T^{*}\mathcal{H}_{+}$ of $\mathcal{H}_{+}$.
Let $\{\phi^{i}\},\{\phi_{i}\}\subset H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}S_{T}$ be bases dual with respect to orbifold Poincaré pairing, i.e. $(\phi^{i},\phi_{j})_{\text{CR}}=\delta^{i}_{j}$. A general point in $\mathcal{H}$ takes the form
$\sum_{a=0}^{\infty}\sum_{i}p_{a,i}\phi^{i}(-z)^{-a-1}+\sum_{b=0}^{\infty}\sum_{j}q_{b}^{j}\phi_{j}z^{b}.$
This defines Darboux coordinates $\{p_{a,i},q_{b}^{j}\}$ on $(\mathcal{H},\Omega)$ compatible with the polarization $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$. Let $p_{a}=\sum_{i}p_{a,i}\phi^{i},q_{b}=\sum_{j}q_{b}^{j}\phi_{j}$, and denote
$\mathbf{p}=\sum_{k=0}^{\infty}p_{k}(-z)^{-k-1},\quad\mathbf{q}=\sum_{k=0}^{\infty}q_{k}z^{k}.$
Givental’s Lagrangian cone $\mathcal{L}_{X}$ is defined to be the graph of the differential of the genus-zero descendant potential $\mathcal{F}^{0}_{X}$,
$\mathcal{L}_{X}=\left\{(\mathbf{p},\mathbf{q})\in\mathcal{H}_{+}\oplus\mathcal{H}_{-}:\mathbf{p}=d_{\mathbf{q}}\mathcal{F}^{0}_{X}\right\}\subset T^{*}\mathcal{H}_{+}\cong\mathcal{H}.$
Let $x=(x_{1},\dots,x_{m})$ be formal variables, an $S_{T}\llbracket Q,x\rrbracket$-valued point of $\mathcal{L}_{X}$ is an element of the form
$1z+\mathbf{t}+\sum_{n=0}^{\infty}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{i}\frac{Q^{d}}{n!}\left\langle\mathbf{t},\dots,\mathbf{t},\frac{\phi_{i}}{-z-\psi}\right\rangle^{X}_{0,n+1,d}\phi^{i}$
for some $\mathbf{t}=\mathbf{t}(z)\in\mathcal{H}_{+}\llbracket x\rrbracket$ satisfying
$\mathbf{t}|_{x=Q=0}=0.$
Let $J$-function be the following formal series
$J_{X}(\mathbf{t},z)=1z+\mathbf{t}+\sum_{n=0}^{\infty}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{i}\frac{Q^{d}}{n!}\left\langle\mathbf{t},\dots,\mathbf{t},\frac{\phi_{i}}{z-\psi}\right\rangle^{X}_{0,n+1,d}\phi^{i}.$
Then $J_{X}(\mathbf{t},-z)$ gives an $S_{T}\llbracket Q,x\rrbracket$-valued point of the Lagrangian cone $\mathcal{L}_{X}$.
Givental’s Lagrangian cone $\mathcal{L}_{X}$ has an important geometric property: each tangent space $\mathcal{T}$ to the cone is tangent to the cone exactly along $z\mathcal{T}$, see [10]. Using the special geometric properties of $\mathcal{L}_{X}$, we can reconstruct the quantum product $\bullet_{\tau}$ from the Lagrangian cone $\mathcal{L}_{X}$ and vice versa.
For instance, let $v\in H^{\bullet}_{\text{CR},T}(X)$, the equivariant quantum connection
$\nabla_{v}:H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}R_{T}[z]\llbracket\tau,Q\rrbracket\rightarrow z^{-1}H^{\bullet}_{\text{CR},T}(X)\otimes_{R_{T}}R_{T}[z]\llbracket\tau,Q\rrbracket$
is defined by
$\nabla_{v}f(\tau)=\partial_{v}f(\tau)+z^{-1}v\bullet_{\tau}f(\tau).$
Denote $\nabla_{\phi_{i}}$ by $\nabla_{i}$. Then the fundamental solution
$L(\tau,z)\in\text{End}_{R_{T}}(H^{\bullet}_{\text{CR},T}(X))\otimes_{R_{T}}R_{T}(\!(z^{-1})\!)\llbracket\tau,Q\rrbracket$
is determined by the following conditions
$\nabla_{i}L(\tau,z)\phi$ $=0,$
$\left(vQ\frac{\partial}{\partial Q}-\partial_{v}\right)L(\tau,z)\phi$ $=L(\tau,z)\frac{v}{z}\phi,$
$L(\tau,z)|_{\tau=Q=0}$ $=\text{id},$
for $i$ runs over the basis $\{\phi_{i}\}$ of $H^{\bullet}_{\text{CR},T}(X)$, $\phi\in H^{\bullet}_{\text{CR},T}(X)$, and $v\in H^{2}_{T}(X)$.
The fundamental solution can be written in the following form using Gromov-Witten invariants:
$L(\tau,z)\phi_{i}=\phi_{i}+\sum_{j}\sum_{d\in\text{NE}(X)_{\mathbb{Z}}}\sum_{n=0}^{\infty}\frac{Q^{d}}{n!}\left\langle\frac{\phi_{i}}{-z-\psi},\tau,\dots,\tau,\phi_{j}\right\rangle^{X}_{0,n+2,d}\phi^{j}.$
The map $\tau\mapsto T_{\tau}=L(\tau,-z)^{-1}\mathcal{H}_{+}$ gives a versal family of tangent spaces to Givental’s Lagrangian cone $\mathcal{L}_{X}$, and the cone can be reconstructed by $\mathcal{L}_{X}=\bigcup_{\tau}zT_{\tau}$.
2. Preliminaries on Toric Geometry
In this Section we present constructions and basic results on toric stacks and toric stack bundles. Our presentation closely follows [8].
2.1. Toric Variety Associated To A Fan
Fix lattices $N,M$ dual to each other and denote $N_{\mathbb{R}}=N\otimes\mathbb{R},M_{\mathbb{R}}=M\otimes\mathbb{R}$. A convex polyhedral cone, or simply a cone, is a set in $N_{\mathbb{R}}$ of the form
$\sigma=\text{Cone}(S)=\left\{\sum_{u\in S}\lambda_{u}u:\lambda_{u}\geq 0\right\},$
where $S\subset N_{\mathbb{R}}$ is finite and call $\sigma$ a cone generated by $S$. The dual cone of $\sigma$ is defined to be
$\sigma^{\vee}=\left\{m\in M_{\mathbb{R}}:\langle m,u\rangle\geq 0\ \text{for all}\ u\in\sigma\right\},$
which is a cone in $M_{\mathbb{R}}$ and $(\sigma^{\vee})^{\vee}=\sigma$. We define the hyperplane
$H_{m}=\left\{u\in N_{\mathbb{R}}:\langle m,u\rangle=0\right\}.$
A face of the cone $\sigma$ is $\tau=H_{m}\cap\sigma$ for some $m\in\sigma^{\vee}$, denoted by $\tau\preceq\sigma$. Every face of $\sigma$ is again a cone. Moreover, an intersection of faces or face of a face is again a face of $\sigma$. A face of codimension $1$ is called a facet and a face of dimension $1$ is called an edge.
A cone is called rational if the generating set $S$ is in $N\subset N_{\mathbb{R}}$. A strongly convex rational cone has minimal generators called the ray generators of the edges. We call such a cone simplicial if its minimal generators are linearly independent over $\mathbb{R}$, and call such a cone smooth if the minimal generators form part of a $\mathbb{Z}$-basis of $N$.
A rational cone $\sigma\subset N_{\mathbb{R}}$ gives a semigroup $S_{\sigma}=\sigma^{\vee}\cap M$, and the affine variety associated to $\sigma$ is
$U_{\sigma}=\text{Spec}(\mathbb{C}[S_{\sigma}]).$
A fan $\Sigma$ in $N_{\mathbb{R}}$ is a finite collection of rational cones $\sigma\subset N_{\mathbb{R}}$ such that for each $\sigma\in\Sigma$, each face of $\sigma$ is also in $\Sigma$, and for any $\sigma_{1},\sigma_{2}\in\Sigma$, the intersection $\sigma_{1}\cap\sigma_{2}$ is a face of each, hence also in $\Sigma$. The support of the fan is $|\Sigma|=\bigcup_{\sigma\in\Sigma}\sigma\subset N_{\mathbb{R}}$.
Given a fan $\Sigma$ in $N_{\mathbb{R}}$, the associated toric variety $X_{\Sigma}$ is defined by gluing the affine toric varieties $U_{\sigma},U_{\sigma^{\prime}}$ along $\sigma\cap\sigma^{\prime}$ for any $\sigma,\sigma^{\prime}\in\Sigma$. In fact, any normal separated toric variety comes from a fan $\Sigma$.
Geometric information of a toric variety can be deduced from the combinatorial data of its fan. For example, $X_{\Sigma}$ is smooth if and only if every cone $\sigma\in\Sigma$ is smooth; $X_{\Sigma}$ is an orbifold if and only if every cone $\sigma\in\Sigma$ is simplicial; $X_{\Sigma}$ is compact if and only if $\Sigma$ is complete, i.e. the support $|\Sigma|=\bigcup_{\sigma\in\Sigma}\sigma$ is the entire $N_{\mathbb{R}}$.
2.1.1. Torus Fixed Points and Toric Blow-ups
There is an orbit-cone correspondence for toric varieties. Let $X_{\Sigma}$ be the toric variety associated to a fan $\Sigma$ in $N_{\mathbb{R}}$. $X_{\Sigma}$ contains the torus $T_{N}=N\otimes\mathbb{C}^{\times}$ as an open subset and $X_{\Sigma}$ has a $T_{N}$-action which extends the action of $T_{N}$ on itself.
Let $\sigma$ be a cone in $\Sigma$, $N_{\sigma}$ the sublattice of $N$ spanned by the points in $\sigma\cap\mathbf{N}$ and $N(\sigma)=N/N_{\sigma}$. The cones $\sigma\in\Sigma$ are in bijective correspondence with $T_{N}$-orbits $O(\sigma)$ of $X_{\Sigma}$. Recall the definition of $U_{\sigma}$, a point $p\in U_{\sigma}$ can be represented by a map $\gamma:S_{\sigma}\rightarrow\mathbb{C}$ which sends $m\in S_{\sigma}$ to $\chi^{m}(p)\in\mathbb{C}$. The torus orbit is
$O(\sigma)$ $=\left\{\gamma:S_{\sigma}\rightarrow\mathbb{C}|\gamma(m)\neq 0\Leftrightarrow m\in\sigma^{\perp}\cap M\right\}\simeq\text{Hom}(\sigma^{\perp}\cap M,\mathbb{C}^{\times})\simeq T_{N(\sigma)}.$
In particular, a top-dimensional cone corresponds to a torus-fixed point of $X_{\Sigma}$.
Operations on a fan $\Sigma$ yield morphisms for the associated toric variety $X_{\Sigma}$. Given a fan $\Sigma$ in $N_{\mathbb{R}}$, a fan $\Sigma^{\prime}$ refines $\Sigma$ if every cone of $\Sigma^{\prime}$ is contained in a cone of $\Sigma$ and $|\Sigma^{\prime}|=|\Sigma|$. Moreover, let $\sigma$ be a smooth top-dimensional cone in $\Sigma$ generated by $u_{1},\dots,u_{n}$, then $u_{1},\dots,u_{n}$ form a basis for $N$. Let $u_{0}=u_{1}+\cdots+u_{n}$ and $\Sigma^{\prime}(\sigma)$ the set of cones generated by subsets of $\{u_{0},u_{1},\dots,u_{n}\}$ not containing $\{u_{1},\dots,u_{n}\}$. Define
$\Sigma^{*}(\sigma)=(\Sigma\setminus\{\sigma\})\cup\Sigma^{\prime}(\sigma),$
the star subdivision of $\Sigma$ along $\sigma$. This is again a fan in $N_{\mathbb{R}}$.
The star subdivision $\Sigma^{*}(\sigma)$ is a refinement of $\Sigma$, and the induced toric morphism
$\phi:X_{\Sigma^{*}(\sigma)}\rightarrow X_{\Sigma}$
is the blow-up of $X_{\Sigma}$ at the distinguished point corresponding to $\sigma$.
Similarly, we can define a toric blow-up along a subvariety $V(\tau)$ using the star subdivision of $\Sigma$ along $\tau$. Let $\tau\in\Sigma$ be such that all cones of $\Sigma$ containing $\tau$ are smooth. Let $u_{\tau}=\sum_{\rho\in\tau(1)}u_{\rho}$ and for each $\sigma\in\Sigma$ containing $\tau$, define
$\Sigma^{*}_{\sigma}(\tau)=\left\{\text{Cone}(A):A\subseteq\{u_{\tau}\}\cup\sigma(1),\tau(1)\nsubseteq A\right\},$
and define the star subdivision of $\Sigma$ along $\tau$ to be
$\Sigma^{*}(\tau)=\left\{\sigma\in\Sigma:\tau\nsubseteq\sigma\right\}\cup\bigcup_{\tau\subseteq\sigma}\Sigma^{*}_{\sigma}(\tau).$
Then the induced toric morphism
$\phi:X_{\Sigma^{*}(\tau)}\rightarrow X_{\Sigma}$
is the blow-up of $X_{\Sigma}$ along $V(\tau)$.
2.2. Toric Stacks
Given $K\cong(\mathbb{C}^{\times})^{r}$, characters of $K$,
$D_{1},\dots,D_{m}\in\mathbb{L}^{\vee}=\text{Hom}(K,\mathbb{C}^{\times}),$
define a map from $K$ to $(\mathbb{C}^{\times})^{m}$, hence an action of $K$ on $\mathbb{C}^{m}$. For a set $I\subset[m]=\{1,2,\dots,m\}$, define
$\angle_{I}$ $=\left\{\sum_{i\in I}a_{i}D_{i}:a_{i}\in\mathbb{R},a_{i}>0\right\}\subset\mathbb{L}^{\vee}\otimes\mathbb{R},$
$(\mathbb{C}^{\times})^{I}\times(\mathbb{C})^{\overline{I}}$ $=\left\{(z_{1},\dots,z_{m}):z_{i}\neq 0\ \text{for}\ i\in I\right\}\subset\mathbb{C}^{m},$
where $\overline{I}=[m]\setminus I$ is the complement of $I$. Set $\angle_{\emptyset}=\{0\}$.
Given a stability condition $\omega\in\mathbb{L}^{\vee}\otimes\mathbb{R}$, we define [^3] Elements of $\mathcal{A}_{\omega}$ are often called anticones.
$\mathcal{A}_{\omega}=\{I\subset[m]:\omega\in\angle_{I}\},\quad U_{\omega}=\bigcup_{I\in\mathcal{A}_{\omega}}(\mathbb{C}^{\times})^{I}\times\mathbb{C}^{\overline{I}},\quad X_{\omega}=[U_{\omega}/K],$
with the assumptions that $m]\in\mathcal{A}_{\omega}$ and for each $I\in\mathcal{A}_{\omega}$, $\{D_{i}:i\in I\}$ spans $\mathbb{L}^{\vee}\otimes\mathbb{R}$ over $\mathbb{R}$.
Let $S\subset[m]$ be the set of indices $i$ such that $m]\setminus\{i\}\notin\mathcal{A}_{\omega}$. Then the characters $\{D_{i}:i\in S\}$ are linearly independent and every $I\in\mathcal{A}_{\omega}$ contains $S$. In fact we may define $X_{\omega}=[U^{\prime}_{\omega}/G]$ where $\mathcal{A}_{\omega}=\{I\sqcup S:I\in\mathcal{A}^{\prime}_{\omega}\}$, $U_{\omega}\cong U^{\prime}_{\omega}\times(\mathbb{C}^{\times})^{|S|}$, $G=\text{Ker}(K\rightarrow(\mathbb{C}^{\times})^{|S|})$ and get the same $X_{\omega}$. And this is the original construction of toric Deligne-Mumford stacks of Borisov-Chen-Smith [3].
The definition here is actually equivalent to the construction by Jiang using an $S$-extended stacky fan [12]. An $S$-extended stacky fan $\mathbf{\Sigma}=(\mathbf{N},\Sigma,\beta,S)$ consists of the following data: a finitely generated abelian group $\mathbf{N}$, a rational simplicial fan $\Sigma$ in $\mathbf{N}\otimes\mathbb{R}$, a homomorphism $\beta:\mathbb{Z}^{m}\rightarrow\mathbf{N}$ with $b_{i}=\beta(e_{i})\in\mathbf{N}$ and $\overline{b_{i}}$ the image of $b_{i}$ in $\mathbf{N}\otimes\mathbb{R}$, and $S\subset[m]$ such that $1$-dimensional cones in $\Sigma$ are in bijective correspondence with $\overline{b_{i}}$ for $i\in[m]\setminus S$, and for each $i\in S$, $\overline{b_{i}}$ lies in the support $\lvert\Sigma\rvert$.
We now explain the equivalence. Starting from the GIT data, consider the following exact sequence
$0\rightarrow\mathbb{L}\overset{(D_{1},\dots,D_{m})}{\xrightarrow{\hskip 28.45274pt}}\mathbb{Z}^{m}\overset{\beta}{\longrightarrow}\mathbf{N}\rightarrow 0,$
i.e. define $\beta:\mathbb{Z}^{m}\rightarrow\mathbf{N}$ to be the cokernel of $\mathbb{L}\rightarrow\mathbb{Z}^{m}$. Let $b_{i}=\beta(e_{i})$ and define $S$ as the set of indices $i$ such that $m]\setminus\{i\}\notin\mathcal{A}_{\omega}$. Then $\Sigma_{\omega}$ is the fan generated by elements $\{\overline{b_{i}},i\in[m]\setminus S\}$ and we get the $S$-extended stacky fan $\mathbf{\Sigma}=(\mathbf{N},\Sigma_{\omega},\beta,S)$.
Conversely, starting from an $S$-extended stacky fan. Using the same exact sequence above, we may define $\mathbb{L}$ as the kernel of $\beta$ and $K=\mathbb{L}\otimes\mathbb{C}^{\times}$. Take the dual exact sequence
$0\rightarrow\mathbf{N}^{\vee}\longrightarrow(\mathbb{Z}^{m})^{\vee}\longrightarrow\mathbb{L}^{\vee},$
and define $D_{i}\in\mathbb{L}^{\vee}$ to be the image of $e_{i}^{\vee}$ in the dual sequence. Define
$\mathcal{A}_{\omega}=\{I\subset[m]:S\subset I\ \text{and}\ \sigma_{\overline{I}}\ \text{is a cone of}\ \Sigma\},$
and take $\omega$ in $\cap_{I\in\mathcal{A}_{\omega}}\angle_{I}$, then this recovers the GIT data.
2.2.1. Equivariant Cohomology of Toric Stacks
Here we summarize some results about equivariant cohomology of toric stacks.
Let $\mathcal{Q}=T/K$, then the $T$-action on $U_{\omega}$ descends to a $\mathcal{Q}$-action on $X_{\omega}=[U_{\omega}/K]$. Let $R_{\mathcal{Q}}=H^{\bullet}_{\mathcal{Q}}(\text{pt};\mathbb{C})$, then $R_{\mathcal{Q}}\cong\text{Sym}^{\bullet}(\mathbf{N}^{\vee}\otimes\mathbb{C})$. Let $\lambda_{i}$ be the equivariant first Chern class of the irreducible $T$-representation given by the projection $T\cong(\mathbb{C}^{\times})^{m}\rightarrow\mathbb{C}^{\times}$ to the $i$th factor, then $R_{T}=\mathbb{C}[\lambda_{1},\dots,\lambda_{m}]$.
The $\mathcal{Q}$-equivariant cohomology ring of $X_{\omega}$ can be written as
$H^{\bullet}_{\mathcal{Q}}(X_{\omega};\mathbb{C})=R_{\mathcal{Q}}[u_{1},\dots,u_{m}]/(\mathfrak{I}+\mathfrak{J}),$
where $u_{i}$ is the $\mathcal{Q}$-equivariant class Poincaré dual to the toric divisor
$\left\{(z_{1},\dots,z_{m})\in U_{\omega}:z_{i}=0\right\}/K],$
and $\mathfrak{I},\mathfrak{J}$ are ideals defined by
$\mathfrak{I}=\left\langle\chi-\sum_{i=1}^{m}\langle\chi,b_{i}\rangle u_{i}:\chi\in\mathbf{N}^{\vee}\otimes\mathbb{C}\right\rangle,\quad\mathfrak{J}=\left\langle\prod_{i\notin I}u_{i}:I\notin\mathcal{A}_{\omega}\right\rangle.$
Note the relations in $\mathfrak{J}$ imply that $u_{i}=0$ in $H^{\bullet}_{\mathcal{Q}}(X_{\omega};\mathbb{C})$ for $i\in S$.
The $T$-equivariant cohomology ring of $X_{\omega}$ is given by the extension of scalars,
$H^{\bullet}_{T}(X_{\omega};\mathbb{C})\cong H^{\bullet}_{\mathcal{Q}}(X_{\omega};\mathbb{C})\otimes_{R_{\mathcal{Q}}}R_{T},$
where the homomorphism $R_{\mathcal{Q}}\rightarrow R_{T}$ is given by $\chi\mapsto\sum_{i=1}^{m}\langle\chi,b_{i}\rangle\lambda_{i}$ for $\chi\in\mathbf{N}^{\vee}\otimes\mathbb{C}$. We also regard $u_{i}$ as a $T$-equivariant class under this identification.
There is a split short exact sequence
(2.1) $0\rightarrow H^{2}_{\mathcal{Q}}(X_{\omega};\mathbb{R})\rightarrow H^{2}_{T}(X_{\omega};\mathbb{R})\rightarrow H^{2}_{K}(\text{pt};\mathbb{R})\rightarrow 0,$
where the splitting is given by $\theta:H^{2}_{K}(\text{pt};\mathbb{R})\cong\mathbb{L}^{\vee}\otimes\mathbb{R}\rightarrow H^{2}_{T}(X_{\omega};\mathbb{R})$, $\theta(D_{i})=u_{i}-\lambda_{i}$. The class $\theta(p)$ actually gives the $T$-equivariant first Chern class of a line bundle $L(p)$ determined by the $K$-representation $p\in\mathbb{L}^{\vee}$.
2.3. Wall and Chamber Structure
The space of stability conditions $\mathbb{L}^{\vee}\otimes\mathbb{R}$ has a wall and chamber structure, which describes how the GIT quotient $X_{\omega}$ changes as $\omega\in\mathbb{L}^{\vee}\otimes\mathbb{R}$ varies.
For each $\omega$, the chamber containing $\omega$ is given by
$C_{\omega}=\bigcap_{I\in\mathcal{A}_{\omega}}\angle_{I}.$
The quotient $X_{\omega^{\prime}}$ is isomorphic to $X_{\omega}$ for any $\omega^{\prime}\in C_{\omega}$. However $X_{\omega^{\prime}}$ changes when $\omega^{\prime}$ crosses a codimension one boundary (called a wall) of $C_{\omega}$.
We also define the cone of ample divisors
$C^{\prime}_{\omega}=\bigcap_{I\in\mathcal{A}^{\prime}_{\omega}}\angle^{\prime}_{I},$
where $\mathcal{A}^{\prime}_{\omega}=\{I\setminus S:I\in\mathcal{A}_{\omega}\}$, $\angle^{\prime}_{I}=\sum_{i\in I}\mathbb{R}_{>0}D^{\prime}_{i}$ and $D^{\prime}_{I}$ is the image of $D_{i}$ in $\mathbb{L}^{\vee}\otimes\mathbb{R}/\sum_{i\in S}\mathbb{R}D_{i}\cong H^{2}(X_{\omega};\mathbb{R})$. Then
$C_{\omega}\cong C^{\prime}_{\omega}\times\left(\sum_{i\in S}\mathbb{R}_{>0}D_{i}\right).$
Let $C_{+}$ and $C_{-}$ be two chambers in $\mathbb{L}^{\vee}\otimes\mathbb{R}$ separated by a hyperplane $W$. Pick stability conditions $\omega_{+}\in C_{+},\omega_{-}\in C_{-}$. Consider $\mathcal{A}_{\pm}=\mathcal{A}_{\omega_{\pm}}$ as defined above, and the corresponding toric stacks $X_{\pm}=X_{\omega_{\pm}}$. Let $\varphi:X_{+}\dashrightarrow X_{-}$ be the birational transformation induced by the toric wall-crossing. We make the additional assumption:
$\sum_{i=1}^{m}D_{i}\in W,$
so that transformation $\varphi$ is crepant. Let $e\in W^{\perp}\cap\mathbb{L}$ be a primitive lattice vector such that $e$ is positive on $C_{+}$ and negative in $C_{-}$.
The chambers in $\mathbb{L}^{\vee}\otimes\mathbb{R}$ form a fan called the Gelfand-Kapranov-Zelevinsky (GKZ) fan. Especially, we consider the subfan consisting of cones $\overline{C_{+}}$ and $\overline{C_{-}}$. Let $\mathcal{M}$ be the toric variety associated to this fan. Since $\overline{C_{\pm}}$ are in general not simplicial cones, $\mathcal{M}$ may be singular. So we may work in some overlattices of $\mathbb{L}$. Define
$\mathbb{K}_{\pm}=\left\{f\in\mathbb{L}\otimes\mathbb{Q}:I_{f}\in\mathcal{A}_{\pm}\right\},$
where $I_{f}=\{i\in[m]:D_{i}\cdot f\in\mathbb{Z}\}$. Let $\tilde{\mathbb{L}}_{\pm}$ be the free $\mathbb{Z}$-submodule generated by $\mathbb{K}_{\pm}$. Note that $D_{j}$ is always in $\tilde{\mathbb{L}}_{\pm}^{\vee}$ for any $j\in S_{\pm}$ and they are linearly independent, so we may extend $\{D_{j}:j\in S_{\pm}\}$ to a basis of $\tilde{\mathbb{L}}_{\pm}^{\vee}$. After reordering, the basis can be written as $\left\{p^{\pm}_{1},\dots,p^{\pm}_{r-1},p^{\pm}_{r}\right\}$ such that $p^{+}_{i}=p^{-}_{i}\in W$ for $1\leq i\leq r-1$ and $p^{\pm}_{r}$ is the unique vector in each basis that does not lie on $W$. Let $\{y^{\pm}_{1},\dots,y^{\pm}_{r}\}$ be the corresponding coordinates, and, for any $d\in\mathbb{L}\otimes\mathbb{R}$, define
$(y^{\pm})^{d}=\prod_{i=1}^{r}(y^{\pm}_{i})^{p^{\pm}_{i}\cdot d}.$
Under change of coordinates, we have $(y^{+})^{e}=(y^{-})^{e}$.
Remark 2.1.
One can interpret $\mathcal{M}$ as the moduli space of certain Landau-Ginzburg models (mirror to the toric stacks). For this paper, our use of $\mathcal{M}$ is mainly to house $y^{\pm}$, which are variables in $I$-functions. See [8] for a detailed discussion about $\mathcal{M}$.
2.4. Toric Stack Bundles
Let $P\rightarrow B$ be a principal $(\mathbb{C}^{\times})^{m}$-bundle over a smooth projective variety $B$. Let $K,D_{i}$ and $U_{\omega}$ be as in Section 2.2. Associated to these data, we define a toric stack bundle to be the quotient
(2.2) $\mathcal{P}={{}^{P}}X_{\omega}=[\left(P\times_{(\mathbb{C}^{\times})^{m}}U_{\omega}\right)/K]\to B,$
where $K$ acts on $U_{\omega}$ via characters $D_{i}$ and acts on $P$ trivially. $\mathcal{P}$ is a fiber bundle with fiber $X_{w}$ over $B$ and it is constructed from the principal $(\mathbb{C}^{\times})^{m}$-bundle $P\to B$ by changing fibers. More details about construction of toric stack bundles can be found in [13].
Let $\mathbf{\Sigma}_{\omega}=(\mathbf{N},\Sigma_{\omega},\beta,S)$ be the corresponding $S$-extended stacky fan corresponding to $K,\{D_{i}\},\omega$ as recalled in Section 2.2, so $X_{\omega}=X(\mathbf{\Sigma}_{\omega})$. The inertia stack of $\mathcal{P}$ is indexed by elements of the following set
$\text{Box}(\mathbf{\Sigma}_{\omega})=\left\{v\in\mathbf{N}:\overline{v}=\sum_{i\notin I}c_{i}\overline{b_{i}}\in\mathbf{N}\otimes\mathbb{R}\ \text{for some}\ I\in\mathcal{A},0\leq c_{i}<1\right\},$
via the decomposition
$I\mathcal{P}=\coprod_{v\in\text{Box}(\mathbf{\Sigma}_{\omega})}\mathcal{P}_{v}=\coprod_{v\in\text{Box}(\mathbf{\Sigma}_{\omega})}{{}^{P}}X(\mathbf{\Sigma}_{\omega}/\sigma(\overline{v})).$
Now define $\mathbb{K}=\left\{f\in\mathbb{L}\otimes\mathbb{Q}:I_{f}\in\mathcal{A}_{\omega}\right\}$ where $I_{f}=\{i\in[m]:D_{i}\cdot f\in\mathbb{Z}\}$. There is a bijection between $\mathbb{K}/\mathbb{L}$ and $\text{Box}(\mathbf{\Sigma}_{\omega})$ given by
$f]\leftrightarrow v_{f}=\sum_{i=1}^{m}\lceil-(D_{i}\cdot f)\rceil b_{i}.$
So we can rewrite the decomposition of $I\mathcal{P}$ over $\mathbb{K}/\mathbb{L}$ as
$I\mathcal{P}=\coprod_{f\in\mathbb{K}/\mathbb{L}}\mathcal{P}_{f}.$
The $T$-equivariant Chen-Ruan cohomology is
$H^{\bullet}_{\text{CR},T}(\mathcal{P})=\bigoplus_{f\in\mathbb{K}/\mathbb{L}}H^{\bullet-2\iota_{f}}_{T}(\mathcal{P}_{f}),$
where $\iota_{f}=\sum_{i\notin I_{f}}\langle D_{i}\cdot f\rangle$ is the age.
Now assume $P$ arises from a direct sum of line bundles $\oplus_{j=1}^{m}L_{j}$ by removing the zero section, and let $\Lambda_{j}=c_{1}(L_{j}^{*})$. We may take $K_{\mathbb{R}}=(S^{1})^{r}$ and regard the quotient (2.2) as a fiberwise symplectic reduction for the action of $K_{\mathbb{R}}$. This defines $r$ tautological bundles over $\mathcal{P}$. Define $P_{1},\dots,-P_{r}$ to be the $T$-equivariant first Chern classes of the tautological bundles, and $p_{1},\dots,p_{r}$ the restrictions of $P_{1},...,P_{r}$ to the fiber $X_{\omega}$. Recall that $\lambda_{1},\dots,\lambda_{m}$ are the first Chern classes of the $m$ duals of tautological bundles on $BT$ and $R_{T}:=H^{\bullet}_{T}(\text{pt};\mathbb{C})=\mathbb{C}[\lambda_{1},\dots,\lambda_{m}]$. The classes $P_{1},\dots,P_{r}$ generate the cohomology ring $H^{\bullet}_{T}(\mathcal{P})$ over $H^{\bullet}(B)\otimes R_{T}$ just as $p_{1},\dots,p_{r}$ generate $H^{\bullet}_{T}(X_{\omega})$ over $R_{T}$. This gives a basis of $H^{\bullet}_{T}(\mathcal{P})$ over $H^{\bullet}(B)\otimes R_{T}$ consisting of monomials in $\{P_{i}\}$.
Recall that we have class $u_{j}$ defined as dual to the $j$-th toric divisor in $X_{\omega}$. Thus we have $u_{j}=\sum_{i=1}^{r}d_{ij}p_{i}$ if we write $D_{j}=(d_{ij})_{i=1}^{r}$. Define
$U_{j}=\sum_{i=1}^{r}d_{ij}P_{i}-\Lambda_{j}-\lambda_{j},$
so $U_{j}$ is the $T$-equivariant cohomology class dual to the $j$-th toric divisor in $\mathcal{P}$. The splitting of the short exact sequence (2.1) gives rise to a map $\theta:H^{2}(B;\mathbb{R})\otimes R_{K}\rightarrow H^{2}_{T}(\mathcal{P};\mathbb{R})$ which maps $D_{j}$ to $U_{j}-\lambda_{j}$ and fixes classes in $H^{2}(B)$. Especially, for a line bundle over $\mathcal{P}$ defined by a character $p$, its $T$-equivariant first Chern class is $\theta(p)$.
2.4.1. Wall-Crossing for Toric Stack Bundles
By the construction of toric stack bundles (2.2), as the stability parameter $\omega$ crosses a wall, $\mathcal{P}$ changes as the fiber $X_{\omega}$ changes.
As in Section 2.3, consider chambers $C_{\pm}\subset\mathbb{L}\otimes\mathbb{R}$ separated by a wall $W$. Let $\omega_{\pm}\in C_{\pm}$ be stability conditions and let
$\mathcal{P}_{\pm}=[P\times_{(\mathbb{C}^{\times})^{m}}U_{\omega_{\pm}}/K]$
[…]
$\tilde{\mathcal{P}}$$\mathcal{P}_{+}$$\mathcal{P}_{-}$$\mathcal{P}_{0},$$\scriptstyle{f_{+}}$$\scriptstyle{f_{-}}$$\scriptstyle{\varphi}$$\scriptstyle{g_{+}}$$\scriptstyle{g_{-}}$
where each map is induced from a map on toric fibers. Especially, the canonical bundles $K_{\mathcal{P}_{\pm}}$ of $\mathcal{P}_{\pm}$ are restrictions of the same line bundle $L_{0}$ on $\mathcal{P}_{0}$ given by a line bundle on $B$ and the character $\sum_{i=1}^{m}D_{i}$. The condition $\sum_{i=1}^{m}D_{i}\in W$ ensures that $L_{0}$ comes from a $\mathbb{Q}$-Cartier divisor $\mathcal{D}_{0}$ on $\mathcal{P}_{0}$. Then we have
$f^{*}_{+}(K_{\mathcal{P}_{+}})=f^{*}_{+}\circ g^{*}_{+}(\mathcal{D}_{0})=f^{*}_{-}\circ g^{*}_{-}(\mathcal{D}_{0})=f^{*}_{-}(K_{\mathcal{P}_{-}}),$
which shows that the birational map $\varphi$ is crepant.
3. Crepant Transformation Correspondence
In this Section we prove Theorem 0.1. First, we compute the analytic continuation of the $I$-function using the Mellin-Barnes method, which gives the linear transformation $\mathbb{U}$. Then we compute the Fourier-Mukai transformation on $K$-groups and compare it with $\mathbb{U}$.
3.1. $I$-Function and $H$-Function
The $I$-function for a toric stack bundle $\mathcal{P}$ is defined in [14] to be
(3.1) $I^{\mathcal{P}}(y,z)=ze^{\sigma/z}\sum_{\mathfrak{D}\in\overline{NE}(B)}J^{B}_{\mathfrak{D}}(z,\tau)Q^{\mathfrak{D}}\sum_{d\in\mathbb{K}}y^{d}\Delta\mathbf{1}_{\left[-d\right]},$
where $J^{B}_{\mathfrak{D}}(z,\tau)$ is the degree-$\mathfrak{D}$ component of the $J$-function of $B$, $\sigma\in H^{2}_{T}(\mathcal{P})$ is a class given by $\sigma=\sum_{i=1}^{r}\theta(p_{i})\text{log}(y_{i})+c_{0}(\lambda)$ where $c_{0}(\lambda)=\lambda_{1}+\cdots+\lambda_{m}$ and $\theta$ is the map defined in Section 2.4, and
$\Delta$ $=\prod_{j=1}^{m}\frac{\prod_{\langle a\rangle=\langle D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\rangle,a\leq 0}(U_{j}+az)}{\prod_{\langle a\rangle=\langle D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\rangle,a\leq D_{j}\cdot d+\Lambda_{j}\mathfrak{D}}(U_{j}+az)}$
$=\prod_{j=1}^{m}\frac{1}{z^{\lceil D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\rceil}}\frac{\Gamma(U_{j}/z+\langle D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\rangle)}{\Gamma(U_{j}/z+D_{j}\cdot d+\Lambda_{j}\mathfrak{D}+1)}.$
The main result of [14] states that $I^{\mathcal{P}}$ lies on the Lagrangian cone for $\mathcal{P}$. Hence $I^{\mathcal{P}}$ can be used to determine genus 0 Gromov-Witten theory of $\mathcal{P}$. See [4] for the case of toric bundles.
Recall the age for $d\in\mathbb{K}$ is defined as $\iota_{d}=\sum_{i\notin I_{d}}\langle D_{i}\cdot d\rangle$, and if $i\in I_{d}$, we have $D_{i}\cdot d\in\mathbb{Z}$ so $\langle D_{i}\cdot d\rangle=0$. So
$\prod_{j=1}^{m}\frac{1}{z^{\lceil D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\rceil}}$ $={z^{-\sum_{j=1}^{m}\lceil D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\rceil}}={z^{-(\sum_{j=1}^{m}\lceil D_{j}\cdot d\rceil+\Lambda_{j}\mathfrak{D})}}$
$={z^{-(\sum_{i=1}^{m}(D_{j}\cdot d+\langle D_{j}\cdot-d\rangle)+(\sum_{i=1}^{m}\Lambda_{j})\mathfrak{D})}}=\frac{1}{z^{\sum_{i=1}^{m}D_{j}\cdot d-c_{1}(P)\mathfrak{D}}}\cdot\frac{1}{z^{\iota_{[-d]}}},$
and the $I$-function can be written as
$I^{\mathcal{P}}$ $=ze^{\sigma/z}\sum_{\mathfrak{D}\in\overline{NE}(B)}J^{B}_{\mathfrak{D}}(z,\tau)Q^{\mathfrak{D}}\sum_{d\in\mathbb{K}}y^{d}\Delta\mathbf{1}_{\left[-d\right]}$
$=ze^{\sigma/z}\sum_{\mathfrak{D}\in\overline{NE}(B)}J^{B}_{\mathfrak{D}}(z,\tau)Q^{\mathfrak{D}}\sum_{d\in\mathbb{K}}\frac{y^{d}}{z^{(D_{1}+\cdots+D_{m})d-c_{1}(P)\mathfrak{D}}}\prod_{j=1}^{m}\frac{\Gamma(1+U_{j}/z-\langle-D_{j}\cdot d\rangle)}{\Gamma(1+U_{j}/z+D_{j}\cdot d+\Lambda_{j}\mathfrak{D})}\frac{\mathbf{1}_{\left[-d\right]}}{z^{\iota_{\left[-d\right]}}}.$
Define the $H$-function by defining its degree-$\mathfrak{D}$ part to be
$H^{\mathfrak{D}}$ $=e^{\sigma/2\pi i}\sum_{d\in\mathbb{K}}y^{d}\prod_{j=1}^{m}\frac{1}{\Gamma(1+U_{j}/2\pi i+D_{j}\cdot d+\Lambda_{j}\mathfrak{D})}\mathbf{1}_{\left[d\right]}.$
Then we have the following relation between the $H$-function and $I$-function:
(3.2) $z^{-1}I^{\mathcal{P}}=\sum_{\mathfrak{D}\in\overline{NE}(B)}J^{B}_{\mathfrak{D}}(z,\tau)Q^{\mathfrak{D}}{z^{\frac{c_{0}(\lambda)}{z}}}z^{c_{1}(P)\mathfrak{D}}z^{-\frac{\text{deg}}{2}}z^{\rho}(\widehat{\Gamma}\cup(2\pi i)^{\frac{\text{deg}_{0}}{2}}\text{inv}^{*}H^{\mathfrak{D}}(z^{\frac{-\text{deg}(y)}{2}}y)),$
where $\widehat{\Gamma}$ is the $T$-equivariant Gamma class given by
$\widehat{\Gamma}=\bigoplus_{f\in\mathbb{K}/\mathbb{L}}\left(\prod_{j=1}^{m}\Gamma(1+U_{j}-\langle D_{j}\cdot f\rangle)\right)\mathbf{1}_{f},$
and $\rho=\theta(D_{1}+\cdots+D_{m})$. The notation deg is the degree operator for Chen-Ruan orbifold cohomology and $\text{deg}_{0}$ is the degree operator for ordinary cohomology, i.e. the degree without age shifting. The degrees for variables are given by $\text{deg}(z)=2$ and
$\sum_{i=1}^{r}(\text{deg}(y_{i}))p_{i}=2\sum_{i=1}^{m}D_{i}.$
By $z^{\frac{-\text{deg}(y)}{2}}y$ we mean
$(z^{\frac{-\text{deg}(y_{1})}{2}}y_{1},\dots,z^{\frac{-\text{deg}(y_{r})}{2}}y_{r}).$
3.2. Analytic Continuation of $H$-function via Localization
For a minimal anticone $\delta\in\mathcal{A}$, there is a corresponding $T$-fixed point in each fiber $X_{\omega}$ over $b\in B$, and together they form a gerbe $B_{\delta}$ over $B$. Let $i_{\delta}:B_{\delta}\rightarrow\mathcal{P}$ be the inclusion map. Similarly, for inertia stack $I\mathcal{P}$, a $T$-fixed gerbe corresponds to a pair $(\delta,f)$ where $\delta\in\mathcal{A}$ and $f\in\mathbb{K}/\mathbb{L}$ satisfy $D_{i}\cdot f\in\mathbb{Z}$ for all $i\in\delta$. Let $i_{(\delta,f)}:B_{(\delta,f)}\rightarrow\mathcal{P}_{f}$ be the inclusion to $I\mathcal{P}$. We have the following restriction of the $H$-function to a fixed gerbe:
$i^{*}_{(\delta,f)}H^{\mathfrak{D}}=\sum_{d\in\mathbb{K},\left[d\right]=f}\frac{y^{d}}{\prod_{j\in\delta}\Gamma(1+D_{j}\cdot d+\Lambda_{j}\mathfrak{D})}\frac{e^{\frac{\sigma(\delta)}{2\pi i}}}{\prod_{j\notin\delta}\Gamma(1+\frac{U_{j}(\delta)}{2\pi i}+D_{j}\cdot d+\Lambda_{j}\mathfrak{D})},$
where we write $\sigma(\delta)=i^{*}_{\delta}\sigma$ and $U_{j}(\delta)=i^{*}_{\delta}U_{j}$ for short.
For the rest of this Section, any notation with index ++ (respectively -) means it is for $\mathcal{P}_{+}$ (respectively $\mathcal{P}_{-}$). Especially, let $I_{\pm}$ denote their $I$-functions and $H^{\mathfrak{D}}_{\pm}$ denote their $H$-functions respectively. $(\delta_{+},f_{+})$ indexes a $T$-fixed gerbe in $I\mathcal{P}_{+}$ and $(\delta_{-},f_{-})$ indexes a $T$-fixed gerbe in $I\mathcal{P}_{-}$. By $\delta_{+}|\delta_{-}$ we mean $\delta_{+}=\{j_{1},\dots,j_{r-1},j_{+}\},\delta_{-}=\{j_{1},\dots,j_{r-1},j_{-}\}$ with $D_{j_{1}},\dots,D_{j_{r-1}}\in W$, $D_{j_{+}}\cdot e>0,D_{j_{-}}\cdot e<0$. By $(\delta_{+},f_{+})|(\delta_{-},f_{-})$ we mean $\delta_{+}|\delta_{-}$ and there exists $\alpha\in\mathbb{Q}$ such that $f_{-}=f_{+}+\alpha e$.
We will show the following relation between $H^{\mathfrak{D}}_{+}$ and $H^{\mathfrak{D}}_{-}$:
Theorem 3.1.
Let $(\delta_{+},f_{+})$ index a T-fixed section on $I\mathcal{P}_{+}$, then for $\delta_{+}\in\mathcal{A}_{+}\cap\mathcal{A}_{-}$, we have
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{-}.$
Otherwise we have
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}i^{*}_{(\delta_{-},f_{-})}H^{\mathfrak{D}}_{-},$
where
$C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}$ $=e^{\frac{\pi iw}{D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}$
$\cdot\frac{\sin\pi(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}{(-D_{j_{-}}\cdot e)\sin\frac{\pi}{-D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}\cdot\prod_{j:D_{j}\cdot e<0,j\neq j_{-}}\frac{\sin\pi(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot f_{+})}{\sin\pi(\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot f_{-})},$
with $w=-1-\sum_{j:D_{j}\cdot e<0}D_{j}\cdot e=-1+\sum_{j:D_{j}\cdot e>0}D_{j}\cdot e$ and $j_{-}\in\delta_{-}$ is the unique element such that $D_{j_{-}}\cdot e<0$. Note that $C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}$ does not depend on $\mathfrak{D}$.
We need the following
Lemma 3.2.
Let $\delta_{-}\in\mathcal{A}_{-}$ and $\delta_{+}\in\mathcal{A}_{+}$ be minimal anticones such that $\delta_{-}|\delta_{+}$, and $j_{-}\in\delta_{-}$ the unique element which is not in $\delta_{+}$. Then, for any $j\in[m]$, we have
$U_{j}(\delta_{+})=U_{j}(\delta_{-})+\frac{D_{j}\cdot e}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+}).$
Proof.
Write $\delta_{-}=\{j_{1},\dots,j_{r-1},j_{-}\}$, then $D_{1},\dots,D_{r-1},D_{j_{-}}$ form a basis of $\mathbb{L}\otimes\mathbb{Q}$, so for any $D_{j}$ we have
(3.3) $D_{j}=c_{1}D_{1}+\cdots+c_{r-1}D_{r-1}+c_{j_{-}}D_{j_{-}}.$
Multiplying (3.3) with $e$ gives $D_{j}\cdot e=c_{j_{-}}D_{j_{-}}\cdot e$, or $c_{j_{-}}=\frac{D_{j}\cdot e}{D_{j_{-}}\cdot e}$.
Now apply $\theta$ to (3.3), we have
$U_{j}-\lambda_{j}=c_{1}(U_{1}-\lambda_{1})+\cdots+c_{r-1}(U_{r-1}-\lambda_{r-1})+c_{j_{-}}(U_{j_{-}}-\lambda_{j_{-}}).$
Applying $i^{*}_{\delta_{-}}$ and $i^{*}_{\delta_{+}}$ gives following equations:
$U_{j}(\delta_{-})-\lambda_{j}=\sum_{i=1}^{r-1}c_{i}(-\lambda_{i})+c_{j_{-}}(-\lambda_{j_{-}}),$
$U_{j}(\delta_{+})-\lambda_{j}=\sum_{i=1}^{r-1}c_{i}(-\lambda_{i})+c_{j_{-}}(U_{j_{-}}(\delta_{+})-\lambda_{j_{-}}).$
Taking their difference, we get $U_{j}(\delta_{+})-U_{j}(\delta_{-})=c_{j_{-}}U_{j_{-}}(\delta_{+})$, which is what we need. ∎
Remark 3.3.
The above proof shows that this lemma can be generalized to $i^{*}_{\delta_{+}}\theta_{+}(p)=i^{*}_{\delta_{-}}\theta_{-}(p)+\frac{p\cdot e}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})$ for any $p\in\mathbb{L}\otimes\mathbb{C}$. Especially, we have $\sigma_{+}(\delta)=\sigma_{-}(\delta)$ for $\delta\in\mathcal{A}_{-}\cap\mathcal{A}_{+}$ and $\sigma_{+}(\delta_{+})=\sigma_{-}(\delta_{-})+\frac{(\text{log}(y^{+}))^{e}}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})$ for $\delta_{+}|\delta_{-}$.
Proof of Theorem 3.1.
For $\delta\in\mathcal{A}_{+}\cap\mathcal{A}_{-}$, the birational map $\varphi:\mathcal{P}_{+}\dashrightarrow\mathcal{P}_{-}$ is an isomorphism in a neighborhood of $B_{\delta}$, and we have $\sigma_{+}(\delta)=\sigma_{-}(\delta)$, so $i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{-}$.
Consider $\delta_{+}\in\mathcal{A}_{+}\setminus\mathcal{A}_{-}$, restricting $H^{\mathfrak{D}}_{+}$ to $(\delta_{+},f_{+})$ gives
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=\sum_{d\in\mathbb{K}_{+},\left[d\right]=f_{+}}\frac{(y^{+})^{d}}{\prod_{j\in\delta_{+}}\Gamma(1+D_{j}\cdot d+\Lambda_{j}\mathfrak{D})}\frac{e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}}{\prod_{j\notin\delta_{+}}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d+\Lambda_{j}\mathfrak{D})}.$
Define $\delta^{\vee}_{+}=\{d\in\mathbb{L}\otimes\mathbb{R}:D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\geq 0,\forall j\in\delta_{+}\}$. Note that the summation in $i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}$ is actually over $d\in\delta^{\vee}_{+}$. Also for $j\in\delta_{+}$, $D_{j}\cdot e=0$ except for one $j_{+}$ where $D_{j_{+}}\cdot e>0$, so for any $d\in\delta^{\vee}_{+}$ we may write $d=d_{+}+ke$ where $d_{+}\in\delta^{\vee}_{+}$ but $d_{+}-e\notin\delta^{\vee}_{+}$, $k\in\mathbb{Z}_{\geq 0}$. Then
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=\sum_{d_{+}\in\delta^{\vee}_{+},\left[d_{+}\right]=f_{+}}(y^{+})^{d_{+}}\sum_{k=0}^{\infty}\frac{e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}((y^{+})^{e})^{k}}{\prod_{j=1}^{m}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}.$
Using the identity $\Gamma(y)\Gamma(1-y)=\pi/\sin(\pi y)$ to rewrite the second sum over $k$, we have:
$\sum_{k=0}^{\infty}\frac{e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}((y^{+})^{e})^{k}}{\prod_{j=1}^{m}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}$
$=$ $\sum_{k=0}^{\infty}e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}((y^{+})^{e})^{k}\prod_{j:D_{j}\cdot e<0}\frac{(-1)^{kD_{j}\cdot e+\Lambda_{j}\mathfrak{D}}\sin\pi(-\frac{U_{j}(\delta_{+})}{2\pi i}-D_{j}\cdot d_{+})}{\pi}$
$\times\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+kD_{j}\cdot e+\Lambda_{j}\mathfrak{D})},$
which is equal to
$\sum_{k=0}^{\infty}e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}\text{Res}_{s=k}\Gamma(s)\Gamma(1-s)e^{\pi is}((y^{+})^{e})^{s}\prod_{j:D_{j}\cdot e<0}\frac{e^{\pi i(sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}\sin\pi(-\frac{U_{j}(\delta_{+})}{2\pi i}-D_{j}\cdot d_{+})}{\pi}$
$\times\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}ds.$
Consider the following integral
$e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}\int_{C}\Gamma(s)\Gamma(1-s)\frac{\prod_{j:D_{j}\cdot e<0}\Gamma(-(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D}))}{\prod_{j:D_{j}\cdot e\geq 0}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}(e^{-\pi i\omega}(y^{+})^{e})^{s}ds.$
It has poles at $s\in\mathbb{Z}$ and
(3.4) $s=\frac{1}{-D_{j_{-}}\cdot e}\left(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}\cdot d_{+}+\Lambda_{j_{-}}\mathfrak{D}-n\right),$
where $D_{j_{-}}\cdot e<0$ and $n$ is a non-negative integer. Choose a contour $C$ such that poles at non-negative integers are on the right of $C$ and other poles are on the left of $C$. In view of [2] and the assumption $\sum_{j}D_{j}\in W$, we can see that this integral is convergent and analytic as a function of $(y^{+})^{e}$ in the domain $\{(y^{+})^{e}:\lvert\text{arg}((y^{+})^{e})-w\pi\rvert<\pi\}$. Moreover, let $\mathfrak{c}=\prod_{j:D_{j}\cdot e\neq 0}(D_{j}\cdot e)^{D_{j}\cdot e}$ be the so-called conifold point, then for $\lvert(y^{+})^{e}\rvert<\lvert\mathfrak{c}\rvert$ the integral is equal to the sum of residues on the right of $C$ and for $\lvert(y^{+})^{e}\rvert>\lvert\mathfrak{c}\rvert$ the integral is equal to the negative of the sum of residues on the left of $C$. By our choice of $C$, for $\lvert(y^{+})^{e}\rvert<\lvert\mathfrak{c}\rvert$ the integral, which is equal to the sum of residues on the right of $C$, gives part of the localization of $H^{\mathfrak{D}}_{+}$, and we define the analytic continuation of $H^{\mathfrak{D}}_{+}$ using the sum of residues on the left of $C$ for $\lvert(y^{+})^{e}\rvert>\lvert\mathfrak{c}\rvert$.
To compute the sum of residues on the left, first note that the residues at poles s=1ns=-1-n vanish, so we only need to consider poles (3.4). By straightforward computations, this is
(3.5) $\begin{split}&-e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}((y^{+})^{e})^{s}e^{\pi i[s(1+D_{j_{-}}\cdot e)+\Lambda_{j_{-}}\mathfrak{D}]}\frac{\sin\pi(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}\cdot d_{+})}{\sin\pi s}\frac{1}{-D_{j_{-}}\cdot e}\frac{(-1)^{n}}{n!}\\ &\times\prod_{j:D_{j}\cdot e<0,j\neq j_{-}}\frac{e^{\pi i(sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}\sin\pi(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+})}{\sin\pi(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}\\ &\times\prod_{j:j\neq j_{-}}\frac{1}{\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+sD_{j}\cdot e+\Lambda_{j}\mathfrak{D})}.\end{split}$
Since $D_{j_{-}}\cdot e<0$, $D_{j_{-}}\cdot e\in\mathbb{Z}_{+}$, we may write the non-negative integer $n$ as $n=k(-D_{j_{-}}e)+\Lambda_{j_{-}}\mathfrak{D}+l$ with $0\leq l+\Lambda_{j_{-}}\mathfrak{D}<(-D_{j_{-}}e)$. Let $d_{-}=d_{+}+\frac{D_{j_{-}}d_{+}-l}{-D_{j_{-}}e}e$, then
$\Lambda_{j_{-}}\mathfrak{D}\leq l=D_{j_{-}}\cdot d_{-}<-D_{j_{-}}e-\Lambda_{j_{-}}\mathfrak{D},$
so we have
$0\leq D_{j_{-}}\cdot d_{-}+\Lambda_{j_{-}}\mathfrak{D}<-D_{j_{-}}\cdot e.$
Recall our definition of $\delta^{\vee}_{+}=\{d\in\mathbb{L}\otimes\mathbb{R}:D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\geq 0,\forall j\in\delta_{+}\}$, we have $\delta^{\vee}_{-}=\{d\in\mathbb{L}\otimes\mathbb{R}:D_{j}\cdot d+\Lambda_{j}\mathfrak{D}\geq 0,\forall j\in\delta_{-}\}$. So $d_{-}\in\delta^{\vee}_{-}$ but $d_{-}+e\notin\delta^{\vee}_{-}$.
By Lemma 3.2, we also have
$\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+s(D_{j}\cdot e)=\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e),$
and
$\sigma_{+}(\delta_{+})=\sigma_{-}(\delta_{-})+\frac{(\text{log}(y^{+}))^{e}}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+}).$
Now we rewrite the residue (3.5) to relate it to $i^{\star}_{(\delta_{-},f_{-})}H^{\mathfrak{D}}_{-}$ using these relations:
(3.6) $\begin{split}&-e^{\frac{\sigma_{-}(\delta_{-})}{2\pi i}}((y^{+})^{d_{-}-d_{+}-ke})e^{\frac{\pi iw}{D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}\cdot(d_{+}-d_{-}))}\frac{\sin\pi(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}\cdot(d_{+}-d_{-}))}{(-D_{j_{-}}\cdot e)\sin\frac{\pi}{-D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}\cdot(d_{+}-d_{-}))}\\ &\times\prod_{j:D_{j}e<0,j\neq j_{-}}\frac{\sin\pi(\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+})}{\sin\pi(\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-})}\prod_{j=1}^{m}\frac{1}{\Gamma(1+\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e)+\Lambda_{j}\mathfrak{D})}.\end{split}$
Let $f_{-}$ be the equivalent class of $d_{-}$ in $\mathbb{K}_{-}/\mathbb{L}$, then we have the relation $d_{+}-d_{-}=f_{+}-f_{-}+Ne$ for some $N\in\mathbb{Z}$. Take this substitution into (3.6) and the dependence on $N$ cancels, so (3.6) is in fact equal to
$e^{\frac{\sigma_{-}(\delta_{-})}{2\pi i}}((y^{+})^{d_{-}-d_{+}-ke})C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}\prod_{j=1}^{m}\frac{1}{\Gamma(1+\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e)+\Lambda_{j}\mathfrak{D})}.$
Now apply the above result to analytic continuation of $i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}$. Note that the summation over residues on the left of $C$ is the same as the summation over $(\delta_{-},f_{-})$ such that $(\delta_{-},f_{-})|(\delta_{+},f_{+})$. We have
$i^{*}_{(\delta_{+},f_{+})}H^{\mathfrak{D}}_{+}=\sum_{d_{+}\in\delta^{\vee}_{+},\left[d_{+}\right]=f_{+}}(y^{+})^{d_{+}}\sum_{k=0}^{\infty}\frac{e^{\frac{\sigma_{+}(\delta_{+})}{2\pi i}}((y^{+})^{e})^{k}}{\prod_{j=1}^{m}\Gamma(1+\frac{U_{j}(\delta_{+})}{2\pi i}+D_{j}\cdot d_{+}+k(D_{j}\cdot e)+\Lambda_{j}\mathfrak{D})}$
$=\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}\sum_{d_{-}\in\delta^{\vee}_{-},\left[d_{-}\right]=f_{-}}(y^{-})^{d_{-}}\sum_{k=0}^{\infty}\frac{e^{\frac{\sigma_{-}(\delta_{-})}{2\pi i}}((y^{-})^{e})^{-k}}{\prod_{j=1}^{m}\Gamma(1+\frac{U_{j}(\delta_{-})}{2\pi i}+D_{j}\cdot d_{-}-k(D_{j}\cdot e)+\Lambda_{j}\mathfrak{D})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}$
$=\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}i^{*}_{(\delta_{-},f_{-})}H^{\mathfrak{D}}_{-}.$
This completes the proof. ∎
Define the following linear transformation: for $\alpha\in H_{\text{CR},T}^{\bullet}(\mathcal{P}_{+})$,
$\mathbb{U}_{H}(\alpha)=\sum_{(\delta,f):\delta\in\mathcal{A}_{+}\cap\mathcal{A}_{-}}(i^{*}_{(\delta,f)}\alpha)\frac{\mathbf{1}_{\delta,f}}{e_{T}(N_{\delta,f})}+\sum_{(\delta_{+},f_{+}):\delta_{+}\in\mathcal{A}_{+}\setminus\mathcal{A}_{-}}\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}(i^{*}_{(\delta_{-},f_{-})}\alpha)\frac{\mathbf{1}_{\delta_{+},f_{+}}}{e_{T}(N_{\delta_{+},f_{+}})}.$
Then it is straightforward to check that $H^{\mathfrak{D}}_{+}=\mathbb{U}_{H}H^{\mathfrak{D}}_{-}$ for all $\mathfrak{D}\in\overline{NE}(B)$.
Define a transformation $\mathbb{U}$ to be the one making the following diagram commute:
(3.7) $H^{\bullet\bullet}_{T}(I\mathcal{P}_{-})\otimes_{\widehat{R}_{T}}\widehat{S}_{T}$$H^{\bullet\bullet}_{T}(I\mathcal{P}_{+})\otimes_{\widehat{R}_{T}}\widehat{S}_{T}$$\tilde{\mathcal{H}}(\mathcal{P}_{-}):=H^{\bullet}_{CR,T}(\mathcal{P}_{-})\otimes_{R_{T}}S_{T}[\text{log}z]((z^{-1/k}))$$H^{\bullet}_{CR,T}(\mathcal{P}_{+})\otimes_{R_{T}}S_{T}[\text{log}z]((z^{-1/k}))=:\tilde{\mathcal{H}}(\mathcal{P}_{+}).$$\scriptstyle{\mathbb{U}_{H}}$$\scriptstyle{\tilde{\Psi}_{-}}$$\scriptstyle{\tilde{\Psi}_{+}}$$\scriptstyle{\mathbb{U}}$
[…]
(3.8) $\tilde{\Psi}_{\pm}(\alpha)=z^{-\frac{\text{deg}}{2}}z^{\rho_{\pm}}({\Gamma(T_{B})}\widehat{\Gamma_{\pm}}\cup(2\pi i)^{\frac{\text{deg}_{0}}{2}}\text{inv}^{*}\alpha)|_{Q^{\mathfrak{D}}\mapsto Q^{\mathfrak{D}}z^{c_{1}(\mathcal{P}_{\pm})\cdot\mathfrak{D}}}.$
A direct calculation shows $I_{+}=\mathbb{U}I_{-}$. This proves part (1) of Theorem 0.1.
Part (2) of Theorem 0.1 also follows immediately. Recall that $p^{+}_{i}=p^{-}_{i},1\leq i\leq r-1$ form a basis for the wall $W$, so part (2) of Theorem 0.1 is equivalent to $\theta_{+}(p^{+}_{i})\circ\mathbb{U}=\mathbb{U}\circ\theta_{-}(p^{-}_{i})$ for $1\leq i\leq r-1$, which follows from definitions of $\theta$ and $\mathbb{U}$.
3.3. Fourier-Mukai Transformation
Consider the following diagram:
$\tilde{\mathcal{P}}$$\mathcal{P}_{+}$$\mathcal{P}_{-},$$\scriptstyle{f_{+}}$$\scriptstyle{f_{-}}$$\scriptstyle{\varphi}$
where $\mathcal{P}_{+},\mathcal{P}_{-}$ are defined as above and $\tilde{\mathcal{P}}$ is the toric stack bundle with fiber $\tilde{X}$ which is the toric blow-up of $X_{+},X_{-}$, and $f_{+},f_{-}$ are induced by the blow-up maps of the fibers. The Fourier-Mukai functor $R(f_{+})_{*}\circ Lf_{-}^{*}:D^{b}_{T}(\mathcal{P}_{-})\to D^{b}_{T}(\mathcal{P}_{+})$ is an equivalence of category, see [5]. We consider the isomorphism $\mathbb{FM}:K^{0}_{T}(\mathcal{P}_{-})\rightarrow K^{0}_{T}(\mathcal{P}_{+})$ induced by this Fourier-Mukai equivalence.
Recall that the fibers $X_{\pm}$ of the toric stack bundle $\mathcal{P}_{\pm}$ are defined from characters $D_{1},\dots,D_{m}\in\mathbb{L}^{\vee}$, stability conditions $\omega_{\pm}\in\mathbb{L}^{\vee}\otimes\mathbb{R}$ and we have the short exact sequence:
$0\rightarrow\mathbb{L}\overset{(D_{1},\dots,D_{m})}{\xrightarrow{\hskip 28.45274pt}}\mathbb{Z}^{m}\overset{\beta}{\longrightarrow}\mathbf{N}\rightarrow 0.$
The toric blow-up $\tilde{X}$ can be defined from $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$ and a stability condition $\tilde{\omega}$. Consider the following short exact sequence:
$0\rightarrow\mathbb{L}\oplus\mathbb{Z}\overset{(\tilde{D}_{1},\dots,\tilde{D}_{m+1})}{\xrightarrow{\hskip 56.9055pt}}\mathbb{Z}^{m+1}\overset{\tilde{\beta}}{\longrightarrow}\mathbf{N}\rightarrow 0,$
where $\tilde{D}_{i}$ is defined by
$\tilde{D_{i}}=\begin{cases}D_{i}\oplus 0&\text{ if }1\leq j\leq m\ \text{and}\ D_{i}\cdot e\leq 0,\\ D_{i}\oplus(-D_{i}\cdot e)&\text{ if }1\leq j\leq m\ \text{and}\ D_{i}\cdot e>0,\\ 0\oplus 1&\text{ if }j=m+1.\end{cases}$
Let $C_{\pm}$ be the chamber containing $\omega_{\pm}$ and $W$ the wall between the two chambers. The stability condition $\tilde{\omega}\in(\mathbb{L}\oplus\mathbb{Z})^{\vee}\otimes\mathbb{R}$ is defined as $(\omega_{0},-\epsilon)$ where $\omega_{0}$ is in the relative interior of $W\cap\overline{C_{+}}=W\cap\overline{C_{-}}$ and $\epsilon$ is a small positive real number. By [8, Lemma 6.16], the toric stack $\tilde{X}$ defined from the above data is the common toric blow-up of $X_{\pm}$.
A calculation of the $K$-theory of toric stack bundles can be found in [13]. We first give a brief description of $K^{0}_{T}(\mathcal{P})$ for a toric stack bundle $\mathcal{P}$. Given an irreducible $K$-representation $p\in\text{Hom}(K,\mathbb{C}^{\times})=\mathbb{L}^{\vee}$, we define a line bundle $L(p)$ over $\mathcal{P}$ as follow:
$L(p)=(P\times_{(\mathbb{C}^{\times})^{m}}U_{\omega})\times\mathbb{C}/(b,z,v)\sim(b,gz,g(p)v),$
where $g\in K,b\in B,z\in U_{\omega},v\in\mathbb{C}$. This line bundle defines a class in $K^{0}_{T}(\mathcal{P})$. Especially, let $p=D_{i}$, we define the following class in $K^{0}_{T}(\mathcal{P}_{\pm})$:
$R^{\pm}_{i}=L_{\pm}(D_{i})\otimes e^{\lambda_{i}},$
where $\pm$ stands for line bundles over $\mathcal{P}_{\pm}$ respectively, $e^{\lambda_{i}}\in\mathbb{C}[T]$ stands for the irreducible $T$-representation given by the $i$-th projection $T\rightarrow\mathbb{C}^{\times}$. Note that these line bundles correspond to the toric divisors, and their inverses are denoted by $S^{\pm}_{i}$.
[…]
Recall that a $T$-fixed locus, which is a gerbe over $B$, is indexed by a minimal anticone $\delta\in\mathcal{A}$. Let $i_{\delta}:B_{\delta}\rightarrow\mathcal{P}$ be the inclusion of this fixed locus. Let $G_{\delta}$ be the common isotropy group of $B_{\delta}$ and $\varrho\in\text{Hom}(G_{\delta},\mathbb{C}^{\times})$ an irreducible representation, then choose a lift $\hat{\varrho}\in\text{Hom}(K,\mathbb{C}^{\times})=\mathbb{L}$. A localized basis for $K^{0}_{T}(\mathcal{P})$ over $K^{0}(B)\otimes\mathbb{C}[T]$ is given by
$\left\{e_{\delta,\varrho}:=(i_{\delta})_{*}\varrho=L(\hat{\varrho})\prod_{i\notin\delta}(1-S_{i})\right\}.$
[…]
Consider a minimal anticone $\tilde{\delta}=\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in\mathcal{A}_{\tilde{\omega}}$ for $\tilde{\mathcal{P}}$ with all $j_{1},\dots,j_{r-1},j_{-},j_{+}\in[m]$, $D_{j_{1}}\cdot e=\cdots=D_{j_{r-1}}\cdot e=0$, $D_{j_{-}}\cdot e<0$ and $D_{j_{+}}\cdot e>0$. Let ${\delta_{-}}=\{j_{1},\dots,j_{r-1},j_{-}\}\in\mathcal{A}_{-}$, ${\delta_{+}}=\{j_{1},\dots,j_{r-1},j_{+}\}\in\mathcal{A}_{+}$ be the corresponding anticones for $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$. Then we have the following diagram:
$B_{\delta_{-}}$$\mathcal{P}_{-}$$B_{\tilde{\delta}}$$\tilde{\mathcal{P}}$$B_{\delta_{+}}$$\mathcal{P}_{+}.$$\scriptstyle{i_{\delta_{-}}}$$\scriptstyle{f_{-,\tilde{\delta}}}$$\scriptstyle{f_{+,\tilde{\delta}}}$$\scriptstyle{i_{\tilde{\delta}}}$$\scriptstyle{f_{-}}$$\scriptstyle{f_{+}}$$\scriptstyle{i_{\delta_{+}}}$
Lemma 3.4.
For a Laurent polynomial $q$, we have
$(f_{+,\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}L(p,n)q(\tilde{R_{1}},\dots,\tilde{R}_{m+1})=(i_{\delta_{+}})^{*}\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(p)t^{p\cdot e+n}q(t^{-l_{1}}R^{+}_{1},\dots,t^{-l_{m}}R^{+}_{m},t),$
where $l=-D_{j_{-}}\cdot e$ and $\mathcal{T}=\{\zeta\cdot(R^{+}_{j_{-}})^{1/l}:\zeta\in\boldsymbol{\mu}_{l}\}$.
Proof.
The common isotropy groups for $B_{\tilde{\delta}}$ and $B_{\delta_{+}}$ are given by
$G_{\tilde{\delta}}$ $=\{(g,\lambda)\in K\times\mathbb{C}^{\times}:D_{j_{-}}(g)=1,D_{j}(g)\lambda^{-D_{j}\cdot e}=1\ \text{for}\ j\in\delta_{+}\},$
$G_{\delta_{+}}$ $=\{h\in K:D_{j}(h)=1\ \text{for all}\ j\in\delta_{+}\}.$
The map $f_{+,\tilde{\delta}}$ is induced by a homomorphism from $G_{\tilde{\delta}}$ to $G_{\delta_{+}}$, $\phi_{+}:(g,\lambda)\mapsto h=g\lambda^{-e}$. The kernel of this map is $\{(\lambda^{e},\lambda):\lambda\in\boldsymbol{\mu}_{l}\}$, here $\boldsymbol{\mu}_{l}$ is the cyclic group of order $l$. This gives the following exact sequence
$1\rightarrow\boldsymbol{\mu}_{l}\rightarrow G_{\tilde{\delta}}\rightarrow G_{\delta_{+}}\rightarrow 1.$
This actually shows that $f_{+,\tilde{\delta}}$ makes $B_{\tilde{\delta}}$ a $\boldsymbol{\mu}_{l}$-gerbe over $B_{\delta_{+}}$. Now given a character $(p,n)\in\text{Hom}(K\times\mathbb{C}^{\times},\mathbb{C}^{\times})$, recall that the line bundle $L(p,n)$ is defined by
$L(p,n)=\left((\tilde{P}\times_{(\mathbb{C}^{\times})^{m+1}}U_{\tilde{\omega}})\times\mathbb{C}\right)/(b,z,v)\sim(b,(g,\lambda)\cdot z,p(g)\lambda^{n}v),$
where $\tilde{P}$ is a principal $(\mathbb{C}^{\times})^{m+1}$-bundle, $U_{\tilde{\omega}}$ is defined by the GIT data $\tilde{\omega}$, $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$ and $\mathcal{A}_{\tilde{\omega}}$ as introduced in Section 2.2, and for any $(g,\lambda)\in K\times\mathbb{C}^{\times}$, $z=(z_{1},\dots,z_{m+1})\in U_{\tilde{\omega}}$, the action is defined by $\tilde{D}_{1},\dots,\tilde{D}_{m+1}$, especially we have
$(g,\lambda)\cdot z=(\tilde{D}_{1}(g,\lambda)z_{1},\dots,\tilde{D}_{m}(g,\lambda)z_{m},\lambda z_{m+1}).$
First we consider $(i_{\tilde{\delta}})^{*}L(p,n)$. On each fiber, the line bundle is pulled back to the fixed point corresponding to $\tilde{\delta}=\tilde{\delta}=\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}$:
$\{(z_{1},\dots,z_{m+1})\in U_{\tilde{\omega}}:z_{i}=0\ \text{if}\ i\notin\tilde{\delta}\}/(K\times\mathbb{C}^{\times}).$
In other words, the possibly non-zero coordinates are $(z_{j_{1}},\dots,z_{j_{r-1}},z_{j_{-}},z_{j_{+}})$, and the action of $(g,\lambda)$ on these coordinates is given by
$(D_{j_{1}}(g)z_{j_{1}},\dots,D_{j_{r-1}}(g)z_{j_{r-1}},D_{j_{-}}(g)z_{j_{-}},D_{j_{+}}(g)\lambda^{-D_{j_{+}}\cdot e}z_{j_{+}}).$
Now take $(g,\lambda)=(\lambda^{e},\lambda),\lambda\in\boldsymbol{\mu}_{l}$ and consider the equivalence relation in $L(p,n)$, we have $(b,z,v)\sim(b,z,\lambda^{p\cdot e+n}v)$ when restricting to the fixed point, so it vanishes if $l$ does not divide $p\cdot e+n$. If $l$ divides $p\cdot e+n$, first we have
$(f_{+,\tilde{\delta}})^{*}(i_{\delta_{+}})^{*}L_{+}(p)(R^{+}_{j_{-}})^{(p\cdot e+n)/l}$ $=(i_{\tilde{\delta}})^{*}(f_{+})^{*}L_{+}(p)(R^{+}_{j_{-}})^{(p\cdot e+n)/l}$
$=(i_{\tilde{\delta}})^{*}L(p,-p\cdot e)(\tilde{R}_{j_{-}})^{(p\cdot e+n)/l}(\tilde{R}_{m+1})^{p\cdot e+n}$
$=(i_{\tilde{\delta}})^{*}L(p,-p\cdot e)(\tilde{R}_{m+1})^{p\cdot e+n}$
$=(i_{\tilde{\delta}})^{*}L(p,n).$
[…]
$(f_{+,\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}L(p,n)=(i_{\delta_{+}})^{*}L_{+}(p)(R^{+}_{j_{-}})^{(p\cdot e+n)/l}.$
We can write the two cases together as
$(f_{+,\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}L(p,n)=(i_{\delta_{+}})^{*}\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(p)t^{p\cdot e+n},$
where $\mathcal{T}=\{\zeta\cdot(R^{+}_{j_{-}})^{1/l}:\zeta\in\boldsymbol{\mu}_{l}\}$: this is because
$\frac{1}{l}\sum_{t\in\mathcal{T}}t^{n}=\left\{\begin{matrix}(R^{+}_{j_{-}})^{n/l}\ &\text{if}\ l\ \text{divides}\ n,\\ 0&\text{otherwise}.\end{matrix}\right$.
To prove the lemma, it suffices to consider the case of monomials $q(\tilde{R}_{1},\dots,\tilde{R}_{m+1})=\prod_{i=1}^{m+1}\tilde{R}_{i}^{r_{i}}$:
$(f_{+,\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}L(p,n)q(\tilde{R_{1}},\dots,\tilde{R}_{m+1})$
$=(f_{+,\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}L(p,n)\prod_{i=1}^{m+1}\tilde{R}_{i}^{r_{i}}$
$=(f_{+,\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}L(p+\sum_{i=1}^{m}r_{i}D_{i},n+r_{m+1}-\sum_{i=1}^{m}r_{i}k_{i})\otimes e^{\sum_{i=1}^{m}r_{i}\lambda_{i}}$
$=(i_{\delta_{+}})^{*}\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(p+\sum_{i=1}^{m}r_{i}D_{i})t^{p\cdot e+n+r_{m+1}-\sum_{i=1}^{m}r_{i}k_{i}}\otimes e^{\sum_{i=1}^{m}r_{i}\lambda_{i}}$
$=(i_{\delta_{+}})^{*}\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(p)t^{p\cdot e+n}\prod_{i=1}^{m}(t^{-k_{i}})^{r_{i}}(R^{+}_{i})^{r_{i}}t^{r_{m+1}}$
$=(i_{\delta_{+}})^{*}\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(p)t^{p\cdot e+n}q(t^{-l_{1}}R^{+}_{1},\dots,t^{-l_{m}}R^{+}_{m},t).$
Theorem 3.5.
Let $\delta\in\mathcal{A}_{-}\cap\mathcal{A}_{+}$ be a common minimal anticone, then we have
$\mathbb{FM}(e_{\delta,\varrho})=e_{\delta,\varrho},$
where $\delta$ is regarded as a minimal anticone in $\mathcal{A}_{-}$ on the left and in $\mathcal{A}_{+}$ on the right. Otherwise, for $\delta_{-}\in\mathcal{A}_{-}\setminus\mathcal{A}_{+}$ we have
[…]
Proof.
For $\delta\in\mathcal{A}_{+}\cap\mathcal{A}_{-}$, the birational map $\varphi:\mathcal{P}_{+}\dashrightarrow\mathcal{P}_{-}$ is an isomorphism in a neighborhood of $B_{\delta}$, so $\mathbb{FM}(e_{\delta,\varrho})=e_{\delta,\varrho}$. Now for $\delta_{-}\in\mathcal{A}_{-}\setminus\mathcal{A}_{+}$, let $\delta_{+}\in\mathcal{A}_{+}\setminus\mathcal{A}_{-}$ be such that $\delta_{+}|\delta_{-}$. Consider a minimal anticone $\tilde{\delta}=\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in\mathcal{A}_{~}$ for $\tilde{\mathcal{P}}$ with all $\{j_{1},\dots,j_{r-1},j_{-},j_{+}\}\in[m]$, $D_{j_{1}}\cdot e=\cdots=D_{j_{r-1}}\cdot e=0$, $D_{j_{-}}\cdot e<0$ and $D_{j_{+}}\cdot e>0$. Let ${\delta_{-}}=\{j_{1},\dots,j_{r-1},j_{-}\}\in\mathcal{A}_{-}$, ${\delta_{+}}=\{j_{1},\dots,j_{r-1},j_{+}\}\in\mathcal{A}_{+}$ be the corresponding anticones for $\mathcal{P}_{-}$ and $\mathcal{P}_{+}$. Recall the following diagram:
$B_{\delta_{-}}$$\mathcal{P}_{-}$$B_{\tilde{\delta}}$$\tilde{\mathcal{P}}$$B_{\delta_{+}}$$\mathcal{P}_{+}.$$\scriptstyle{i_{\delta_{-}}}$$\scriptstyle{f_{-,\tilde{\delta}}}$$\scriptstyle{f_{+,\tilde{\delta}}}$$\scriptstyle{i_{\tilde{\delta}}}$$\scriptstyle{f_{-}}$$\scriptstyle{f_{+}}$$\scriptstyle{i_{\delta_{+}}}$
Take $e_{\delta_{-},\varrho}\in K^{0}_{T}(\mathcal{P}_{-})$, then we have $(f_{-})^{*}e_{\delta_{-},\varrho}=(f_{-})^{*}L_{-}(\hat{\varrho})\prod_{i\notin\delta_{-}}(1-S^{-}_{i})=L(\hat{\varrho},0)\prod_{i\notin\delta_{-}}(1-\tilde{S}_{i}\tilde{S}_{m+1}^{k_{i}})$. By localization formula, we have
$(f_{+})_{*}(f_{-})^{*}e_{\delta_{-},\varrho}$ $=(f_{+})_{*}L(\hat{\varrho},0)\prod_{i\notin\delta_{-}}(1-\tilde{S}_{i}\tilde{S}_{m+1}^{k_{i}})$
$=(f_{+})_{*}\sum_{\tilde{\delta}:\delta_{-}\subset\tilde{\delta}}(i_{\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}\frac{L(\hat{\varrho},0)\prod_{i\notin\delta_{-}}(1-\tilde{S}_{i}\tilde{S}_{m+1}^{k_{i}})}{(1-\tilde{S}_{m+1})(\prod_{i\notin\tilde{\delta}}1-\tilde{S}_{i})}$
$=\sum_{\tilde{\delta}:\delta_{-}\subset\tilde{\delta}}(i_{\delta_{+}})_{*}(f_{+,\tilde{\delta}})_{*}(i_{\tilde{\delta}})^{*}\frac{L(\hat{\varrho},0)\prod_{i\notin\delta_{-}}(1-\tilde{S}_{i}\tilde{S}_{m+1}^{k_{i}})}{(1-\tilde{S}_{m+1})(\prod_{i\notin\tilde{\delta}}1-\tilde{S}_{i})}.$
Applying Lemma 3.4, we have
[…]
$=\sum_{\delta_{+}:\delta_{+}|\delta_{-}}(i_{\delta_{+}})_{*}(i_{\delta_{+}})^{*}\frac{\frac{1}{l}\sum_{t\in\mathcal{T}}L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\frac{1-S^{+}_{j_{-}}}{1-t^{-1}}\prod_{i\notin\delta_{-}}(1-t^{-k_{i}}S^{+}_{i})}{\prod_{i\notin\delta_{+}}(1-S^{+}_{i})}$.
We need to check that restriction of
[…]
is only non-zero for fixed points corresponding to minimal anticones $\delta_{+}$ next to $\delta_{-}$.
[…]
$\frac{1-S^{+}_{j_{-}}}{1-t^{-1}}(1-t^{-D_{j_{+}}\cdot e}S^{+}_{j_{+}})=(1-S^{+}_{j_{-}})\frac{1-(t^{-1})^{D_{j_{+}}\cdot e}}{1-t^{-1}}.$
Since $D_{j_{+}}\cdot e$ is a positive integer, this factor is a polynomial in $t^{-1}$ and its restriction to $\delta$ vanishes. So the only possibility for non-zero restriction occurs for $\delta_{+}$ with $\delta_{+}|\delta_{-}$. Applying localization formula one more time, we obtain
[…]
3.4. Comparison
We claim that the Fourier-Mukai transformation is compatible with the analytic continuation above, i.e. the following diagram commutes
(3.9) $K^{0}_{T}(\mathcal{P}_{-})$$K^{0}_{T}(\mathcal{P}_{+})$$H^{\bullet\bullet}_{T}(I\mathcal{P}_{-})$$H^{\bullet\bullet}_{T}(I\mathcal{P}_{+}).$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\tilde{\text{ch}}}$$\scriptstyle{\tilde{\text{ch}}}$$\scriptstyle{\mathbb{U}_{H}}$
Here $\tilde{\text{ch}}$ is the orbifold Chern character.
First, we check commutativity on basis $K^{0}_{T}(B)$. Note that $\tilde{\text{ch}}$ maps $K^{0}_{T}(B)$ to $H^{\bullet}(B)\otimes R_{T}$ and everything is fixed under $\mathbb{FM}$ and $\mathbb{U}_{H}$ in this case.
Now for a basis element $e_{\delta_{-},\varrho}\in K^{0}_{T}(\mathcal{P}_{-})$, if $\delta_{-}\in\mathcal{A}_{-}\cap\mathcal{A}_{+}$, we also have
$\tilde{\text{ch}}(\mathbb{FM}(e_{\delta_{-},\varrho}))=\tilde{\text{ch}}(e_{\delta_{-},\varrho})=\mathbb{U}_{H}\tilde{\text{ch}}(e_{\delta_{-},\varrho}).$
For $\delta_{-}\notin\mathcal{A}_{-}\cap\mathcal{A}_{+}$, we calculate them explicitly. First, note that we have
$)\displaystyle\tilde{\text{ch}}(L_{\pm}(\hat{\varrho}$)) $=\bigoplus_{f\in\mathbb{K}_{\pm}/\mathbb{L}}e^{2\pi i\hat{\varrho}\cdot f}e^{\theta_{\pm}(\hat{\varrho})}\mathbf{1}_{f}.$
Especially we have
$\tilde{\text{ch}}(S^{\pm}_{j})$ $=\bigoplus_{f\in\mathbb{K}_{\pm}/\mathbb{L}}e^{-2\pi iD_{j}\cdot f}e^{-U_{j}}\mathbf{1}_{f},$
and
$\tilde{\text{ch}}(t)$ $=\bigoplus_{f\in\mathbb{K}_{\pm}/\mathbb{L}}\zeta e^{2\pi iD_{j_{-}}\cdot f/l}e^{U_{j_{-}}/l}\mathbf{1}_{f},$
for $t=\zeta(R^{+}_{j_{-}})^{1/l}\in\mathcal{T}$. We may rewrite $\mathbb{FM}(e_{\delta_{-},\varrho})$ as
[…]
It suffices to check the equality locally. Note that $\tilde{\text{ch}}(\mathbb{FM}(e_{\delta_{-},\varrho}))$ is only supported on fixed loci corresponding to $(\delta_{+},f_{+})$ where $\delta_{+}|\delta_{-}$. So we need to check
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}(\mathbb{FM}(e_{\delta_{-},\varrho}))=\sum_{(\delta^{\prime}_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta^{\prime}_{-},f_{-}}_{\delta_{+},f_{+}}(i^{\star}_{(\delta^{\prime}_{-},f_{-})}\tilde{\text{ch}}(e_{\delta_{-},\varrho})).$
First, using Lemma 3.2 we have
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}(L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e})$ $=e^{2\pi i\hat{\varrho}\cdot f_{+}}e^{i_{\delta_{+}}^{\star}\theta_{+}(\hat{\varrho})}e^{-2\pi iD_{j_{-}}\cdot f_{-}\frac{\hat{\varrho}\cdot e}{l}}e^{2\pi iD_{j_{-}}\cdot f_{+}\frac{\hat{\varrho}\cdot e}{l}}e^{U_{j_{-}}(\delta_{+})\frac{\hat{\varrho}\cdot e}{l}}$
$=e^{2\pi i\hat{\varrho}\cdot f_{+}}e^{2\pi i\frac{\hat{\varrho}\cdot e}{l}D_{j_{-}}\cdot(f_{+}-f_{-})}e^{i^{\star}_{\delta_{+}}\theta_{+}(\hat{\varrho})+\frac{\hat{\varrho}\cdot e}{l}U_{j_{-}}(\delta_{+})}$
$=e^{2\pi i(\hat{\varrho}\cdot f_{+}-\alpha\hat{\varrho}\cdot e)}e^{i^{\star}_{\delta_{-}}\theta_{-}(\hat{\varrho})}$
$=e^{2\pi i\hat{\varrho}\cdot f_{-}}e^{i^{\star}_{\delta_{-}}\theta_{-}(\hat{\varrho})}$
$),\displaystyle=i^{\star}_{(\delta_{-},f_{-})}\tilde{\text{ch}}(L_{-}(\hat{\varrho}$)),
where we set $f_{+}=f_{-}+\alpha e$ and $t=e^{-2\pi iD_{j_{-}}\cdot f_{-}/l}(R^{+}_{j_{-}})^{1/l}$. Also
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}(S^{+}_{j}t^{-D_{j}\cdot e})$ $=e^{-2\pi iD_{j}\cdot f_{+}}e^{-U_{j}(\delta_{+})}e^{-2\pi iD_{j_{-}}\cdot f_{-}\frac{-D_{j}\cdot e}{l}}e^{2\pi iD_{j_{-}}\cdot f_{+}\frac{-D_{j}\cdot e}{l}}e^{U_{j_{-}}(\delta_{+})\frac{-D_{j}\cdot e}{l}}$
$=e^{-2\pi i\left(D_{j}\cdot f_{+}-\frac{D_{j}\cdot e}{D_{j_{-}}\cdot e}D_{j_{-}}(f_{+}-f_{-})\right)}e^{-\left(U_{j}(\delta_{+})-\frac{D_{j}\cdot e}{D_{j_{-}}\cdot e}U_{j_{-}}(\delta_{+})\right)}$
$=e^{-2\pi iD_{j}\cdot f_{-}}e^{-U_{j}(\delta_{-})}$
$=i^{\star}_{(\delta_{-},f_{-})}\tilde{\text{ch}}(S_{j}^{-}).$
Together we have
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}\left(L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\prod_{i\notin\delta_{-}}(1-S^{+}_{i}t^{-D_{i}\cdot e})\right)=i^{\star}_{(\delta_{-},f_{-})}\tilde{\text{ch}}(e_{\delta_{-},\varrho}).$
On the other hand, we have
[…]
$=\frac{1}{l}\frac{1-e^{-U_{j_{-}}(\delta_{+})-2\pi iD_{j_{-}}\cdot f_{+}}}{1-e^{-\frac{1}{l}(U_{j_{-}}(\delta_{+})+2\pi iD_{j_{-}}(f_{+}-f_{-}))}}\prod_{i\notin\delta_{-},D_{i}\cdot e<0}\frac{1-e^{-U_{i}(\delta_{+})-2\pi iD_{i}\cdot f_{+}}}{1-e^{-U_{i}(\delta_{-})-2\pi iD_{i}\cdot f_{-}}}$
$=\frac{1}{l}\frac{\sin\left(\frac{-U_{j_{-}}(\delta_{+})}{2i}-\pi D_{j_{-}}\cdot f_{+}\right)}{\sin\left(-\frac{1}{l}(\frac{U_{j_{-}}(\delta_{+})}{2i}+\pi D_{j_{-}}(f_{+}-f_{-}))\right)}\prod_{i\notin\delta_{-},D_{i}\cdot e<0}\frac{\sin\left(\frac{-U_{i}(\delta_{+})}{2i}-\pi D_{i}\cdot f_{+}\right)}{\sin\left(\frac{-U_{i}(\delta_{-})}{2i}-\pi D_{i}\cdot f_{-}\right)}$
$\times e^{-U_{j_{-}}(\delta_{+})/2-\pi iD_{j_{-}}\cdot f_{+}+\frac{1}{2l}(U_{j_{-}}(\delta_{+})+2\pi iD_{j_{-}}(f_{+}-f_{-}))+\sum_{i\notin\delta_{-},D_{i}\cdot e<0}(-U_{i}(\delta_{+})/2-\pi iD_{i}\cdot f_{+}+U_{i}(\delta_{-})/2+\pi iD_{i}\cdot f_{-})}$
$=\frac{1}{l}e^{\frac{\pi iw}{D_{j_{-}}\cdot e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}\frac{\sin\pi(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}{\sin\frac{\pi}{-D_{j_{-}}e}(\frac{U_{j_{-}}(\delta_{+})}{2\pi i}+D_{j_{-}}(f_{+}-f_{-}))}\prod_{i\notin\delta_{-},D_{i}\cdot e<0}\frac{\sin\pi(\frac{U_{i}(\delta_{+})}{2\pi i}+D_{i}f_{+})}{\sin\pi(\frac{U_{i}(\delta_{-})}{2\pi i}+D_{i}f_{-})}$
$=C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}.$
So for any basis element $e_{\delta_{-},\varrho}\in\mathcal{A}_{-}\setminus\mathcal{A}_{+}$ and any $\delta_{+}|\delta_{-}$, we have
[…]
$=\sum_{(\delta_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta_{-},f_{-}}_{\delta_{+},f_{+}}i^{\star}_{(\delta_{-},f_{-})}\tilde{\text{ch}}(e_{\delta_{-},\varrho})$
$=i^{\star}_{(\delta_{+},f_{+})}\mathbb{U}_{H}\tilde{\text{ch}}(e_{\delta_{-},\varrho}),$
since as $f_{-}$ runs over $\mathbb{K}_{-}/\mathbb{L}$, $\zeta=e^{-2\pi iD_{j_{-}}\cdot f_{-}/l}$ runs over $\boldsymbol{\mu}_{l}$. This completes the proof of commutativity of (3.9).
Combining diagrams (3.9) and (3.7), we arrive at part (3) of Theorem 0.1.
The claim that $\mathbb{U}$ is symplectic follows from the commutativity of (0.1), the fact that $\mathbb{FM}$ preserves Mukai pairing, and Toen’s Grothendieck-Riemann-Roch formula.